### v1

# Semantic Equitable Clustering: A Simple, Fast and Effective Strategy for Vision Transformer

Abstract
The Vision Transformer (ViT) has gained prominence for its superior relational modeling prowess. However, its global attention mechanism’s quadratic complexity poses substantial computational burdens. A common remedy spatially groups tokens for self-attention, reducing computational requirements. Nonetheless, this strategy neglects semantic information in tokens, possibly scattering semantically-linked tokens across distinct groups, thus compromising the efficacy of self-attention intended for modeling inter-token dependencies. Motivated by these insights, we introduce a fast and balanced clustering method, named Semantic Equitable Clustering (SEC). SEC clusters tokens based on their global semantic relevance in an efficient, straightforward manner. In contrast to traditional clustering methods requiring multiple iterations, our method achieves token clustering in a single pass. Additionally, SEC regulates the number of tokens per cluster, ensuring a balanced distribution for effective parallel processing on current computational platforms without necessitating further optimization. Capitalizing on SEC, we propose a versatile vision backbone, SecViT. Comprehensive experiments in image classification, object detection, instance segmentation, and semantic segmentation validate to the effectiveness of SecViT. Remarkably, SecViT attains an impressive 84.2% image classification accuracy with only 27M parameters and 4.4G FLOPs, without the need for for additional supervision or data. Code will be available at https://github.com/qhfan/SecViT.
1 Introduction
Since its inception, the Vision Transformer (ViT)[13] has drawn considerable interest from the research community due to its robust modeling prowess. However, the quadratic complexity of Self-Attention leads to significant computational overhead, thus constraining the practicality of ViT. A variety of strategies have been devised to alleviate this computational load, the most prevalent of which involves token grouping, thereby constraining the attention span of each token[36, 12, 53, 48].
Specifically, the Swin-Transformer [36] partitions tokens into multiple small windows, restricting token attention within each window. The CSWin-Transformer [12] adopts a cross-shaped grouping, endowing each token with a global receptive field. MaxViT [48] amalgamates window and grid attention, facilitating intra-window tokens to attend to their counterparts in other windows. However, these methods, solely reliant on spatial positioning, neglect token semantics, potentially restricting the self-attention’s capacity to model semantic dependencies. To mitigate this, DGT [35] employs k-means clustering for query grouping, considering the semantic information of tokens for enhanced feature learning. Nonetheless, the iterative nature of k-means clustering and the potential for uneven token counts per cluster can impact the efficiency of parallel attention operations.
Figure 1: Comparison among Window Partition, Dynamic Group by k-means, and Semantic Equitable Clustering. Our Semantic Equitable Clustering incorporates image semantics while maintaining efficient clustering, eliminating the need for iterative processes such as in k-means. Furthermore, it enables equi-partitioning of tokens, promoting efficient GPU processing without necessitating additional CUDA optimization.
Given these considerations, an optimal token partitioning scheme should efficiently segregate tokens, incorporate semantic information, and efficiently utilize computational resources (e.g., GPU). In response, we introduce a simple, fast and equitable clustering approach named Semantic Equitable Clustering (SEC). SEC segments tokens based on their relevance to global semantic information. Specifically, we employ global pooling to generate a global token encapsulating global semantic information. The similarity between this global token and all other tokens is then computed, reflecting global semantic relevance. Upon obtaining the similarity matrix, tokens (excluding the global token) are sorted by similarity scores, and the tokens with similar scores are grouped into clusters, ensuring uniform token distribution across clusters. As depicted in Fig. 1, SEC comprehensively considers token semantics and completes the clustering process in a single iteration, unlike the multi-iteration k-means. The resulting clusters, containing an equal number of tokens, can be processed in parallel by the GPU efficiently.
Figure 2: Top-1 accuracy v.s. FLOPs on ImageNet-1K of resent SOTA models. SecViT outperforms all the counterparts in all settings.
[…]
2 Related works
Vision Transformer.
The Vision Transformer (ViT) [13] is considered a powerful visual architecture due to its excellent visual modeling capabilities. Many works have improved the Vision Transformer, including enhancing its training efficiency and reducing its computational cost [36, 12, 47, 26, 62]. DeiT uses distillation loss and incorporates extensive data augmentation methods into the ViT training process. This increases the model’s data utilization efficiency, allowing it to train an effective ViT with less data. Hierarchical structures represented by PVT [51, 52, 45, 16, 54] reduce the number of tokens in global attention by downsampling the keys and values (KV), thereby keeping the computational cost of global attention within an acceptable range. In addition to using downsampling to reduce the number of tokens, some methods directly prune tokens based on their importance, retaining important tokens while discarding or merging unimportant ones [15, 42]. This reduces the number of tokens and subsequently lowers the computational cost of the model. Another highly representative approach is to group all tokens such that each token can only attend to tokens within its own group [36, 12, 62, 11, 35]. This method also significantly reduces the computational cost of self-attention.
Grouping-Based Vision Transformer.
Most grouping-based attention mechanisms perform grouping based on spatial structure [36, 12, 48, 11, 35]. Specifically, the Swin-Transformer [36] divides all tokens into equally sized windows based on their spatial positions, where each token can only attend to tokens within its own window. This significantly reduces the model’s computational cost. In contrast, CSWin [12] divides tokens into a crisscrossing structure, which also reduces the model’s computational cost. In addition to dividing tokens into small windows along the spatial dimension, DaViT [11] also splits channels into multiple groups along the channel dimension. Unlike the above methods that only consider positional information for grouping, DGT [35] takes semantic information into account by using k-means clustering to group the queries.
3 Method
3.1 Overall Architecture
The overall architecture of SecViT is shown in Fig. 3(a). For each input image $x\in\mathbb{R}^{3\times H\times W}$, we first process it by a patch embedding to downsample it. Then we obtain the tokens with a shape of $C_{1}\times\frac{H}{4}\times\frac{W}{4}$. Similar to previous work [36, 51, 12], we adopt a hierarchical structure to process tokens. Our model consists of four stages with downsampling factors of $\frac{1}{4}$, $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$, respectively. This structural design facilitates downstream tasks, such as object detection, in constructing feature pyramids.
A SecViT block is composed of three modules: Conditional Positional Encoding (CPE) [7], Semantic Equitable Clustering based Self-Attention (SECA), and classical Feed-Forward Network (FFN) [49]. A complete SecViT block can be represented by Eq. 1:
$\centering\begin{split}&X={\rm CPE(}X_{in}{\rm)}+X_{in},\\ &Y={\rm SECA(}{\rm LN(}X{\rm))}+X,\\ &Z={\rm FFN(}{\rm LN(}Y{\rm))}+Y.\end{split}\@add@centering$ (1)
For each block, the input tensor $X_{in}\in\mathbb{R}^{C\times H\times W}$ is fed into the CPE to introduce the positional information. Then, The Self-Attention based on the Semantic Equitable Clustering (SEC) is employed to serve as the token mixer. The final FFN is utilized to integrate channel-wise information of tokens.
3.2 Semantic Equitable Clustering
As previously mentioned, the design objectives of Semantic Equitable Clustering are threefold: 1) Fully consider the semantic information contained in different tokens during clustering. 2) Unlike k-means and other clustering methods that require multiple iterations, Semantic Equitable Clustering can complete clustering in a single step. 3) Ensure an equal number of tokens in each cluster to facilitate parallel processing on GPUs. In the following paragraphs, we will describe in detail how our Semantic Equitable Clustering achieves these three objectives. And the whole process is illustrated in the Fig. 3(c).
Figure 3: (a) The overall architecture of our Semantic Equitable Clustering Vision Transformer. (b) The composition of each SecViT block. (c) Illustration of our Semantic Equitable Clustering. SEC determines the clustering rules for tokens by calculating the cosine similarity between all tokens and the global token.
Single Clustering Center Related to Semantics.
K-means is relatively complex for two reasons. First, it has multiple cluster centers, and each token needs to calculate its distance to each cluster center to determine its cluster membership. Second, the determination of each cluster center in K-means is not precise and requires multiple iterations to accurately establish the cluster centers.
To address these two issues, we first discard the use of multiple cluster centers and instead calculate the distance between each token and a single center. Based on each token’s distance to this center, we divide the tokens into different intervals. Then, to ensure that our chosen center contains the most comprehensive semantic information, we directly use the result of average pooling of all tokens as the center token. This is because, in most vision foundation models, the output of the average pool is assumed to contain the richest semantic information and is thus used for classification [36, 12, 7, 14].
ToMe [2] indicates that the $K$ contains the richest information from each token. Therefore, we choose the $K$ as the basis for clustering. Specifically, the process for determining the cluster center is shown in Eq. 2:
$\centering\begin{split}Q=W_{Q}X,&K=W_{K}X,V=W_{V}X,\\ &k_{c}={\rm Avg}(K).\end{split}\@add@centering$ (2)
Where $W_{K}$ is a learnable matrix. $k_{c}$ is the determined cluster center. $X$ is the set of input tokens.
Distance Metric Suitable for ViT.
Unlike the Euclidean distance calculation used in the K-means algorithm for computing the distance between tokens, during the actual computation of Self-Attention, similarity between query and key is computed through dot product. To better adapt to the characteristics of Self-Attention, we also measure the distance between tokens using a method similar to dot product. Specifically, we calculate the cosine similarity between the cluster center and each token, and then sort the tokens according to the magnitude of the computed results. The specific process is shown in Eq. 3:
$\centering\begin{split}&sim=\frac{K\cdot k_{c}}{||K||\cdot||k_{c}||},\\ &idx={\rm argsort}(sim),\\ Q^{*}=Q[idx]&,K^{*}=K[idx],V^{*}=V[idx].\end{split}\@add@centering$ (3)
Where $sim$ is the similarity matrix between $K$ and $k_{c}$, the ${\rm argsort}(sim)$ returns the indices of $sim$ sorted in descending order. $Q^{*},K^{*},V^{*}$ are $Q,K,V$ rearranged according to ${\rm argsort}(sim)$.
Equally Partition Tokens based on Distance.
The obtained $Q^{*}$, $K^{*}$, and $V^{*}$ from the previous step have been sorted based on their distances to the cluster center. At this point, we directly group them, so tokens with similar distances to the cluster center are classified into the same cluster. This allows us to directly control an equal number of tokens in each cluster. This process can be clearly illustrated in Fig. 3(c) and denoted as follows:
$\centering\begin{split}Q_{c}=Q_{c\times N_{c}:\left(c+1\right)N_{c}}^{*},K_{c}=K_{c\times N_{c}:\left(c+1\right)N_{c}}^{*},K_{c}=K_{c\times N_{c}:\left(c+1\right)N_{c}}^{*}.\end{split}\@add@centering$ (4)
where $N_{c}$ is the basic token number of each cluster for equal partition and $c$ is the index of the cluster
Based on the above steps, we have completed the clustering process that captures semantic information in the image with minimal sorting cost. Moreover, compared to K-means, we have achieved equi-partitioning of each cluster. After clustering is completed, we apply standard Self-Attention to the tokens within each cluster (SECA), thereby completing the interaction of information between tokens:
$\centering Y_{c}=\text{Attn}\left(Q_{c},K_{c},V_{c}\right).\@add@centering$ (5)
After the completion of SECA, to further enhance the model’s ability to capture local information, we employ a local context enhancement module to model local information, where LCE [62, 12] is a simple depth-wise convolution:
$\centering\begin{split}X={\rm SECA}(X)+{\rm LCE}(W_{V}X),\end{split}\@add@centering$ (6)
4 Experiments
We conducted experiments on a wide range of vision tasks, including image classification on ImageNet-1k [10], object detection and instance segmentation on COCO [33], and semantic segmentation on ADE20K [61]. All models can be trained with 8 A100 80G GPUs.
4.1 ImageNet Classification
Settings.
We train our models from scratch on ImageNet-1k [10]. For a fair comparison, we follow the training strategy in DeiT [47], with classification loss serving as the only supervision. The maximum rates for increasing stochastic depth [23] are set to 0.1, 0.15, 0.4, and 0.5 for SecViT-T, SecViT-S, SecViT-B, and SecViT-L, respectively.
Cost Model Parmas (M) FLOPs (G) Top1-acc (%)
tiny model $\sim 2.5$G PVTv2-b1 [52] 13 2.1 78.7
QuadTree-B-b1 [46] 14 2.3 80.0
RegionViT-T [5] 14 2.4 80.4
MPViT-XS [28] 11 2.9 80.9
VAN-B1 [17] 14 2.5 81.1
BiFormer-T [62] 13 2.2 81.4
Conv2Former-N [22] 15 2.2 81.5
CrossFormer-T [53] 28 2.9 81.5
NAT-M [18] 20 2.7 81.8
FAT-B2 [14] 14 2.0 81.9
QnA-T [1] 16 2.5 82.0
GC-ViT-XT [20] 20 2.6 82.0
SMT-T [34] 12 2.4 82.2
SecViT-T 15 2.5 82.6
small model $\sim 4.5$G Swin-T [36] 29 4.5 81.3
CrossViT-15 [4] 27 5.8 81.5
RVT-S [38] 23 4.7 81.9
ConvNeXt-T [37] 29 4.5 82.1
Focal-T [57] 29 4.9 82.2
MPViT-S [28] 23 4.7 83.0
SG-Former-S [15] 23 4.8 83.2
Ortho-S [25] 24 4.5 83.4
InternImage-T [50] 30 5.0 83.5
GC-ViT-T [20] 28 4.7 83.5
CMT-S [16] 25 4.0 83.5
MaxViT-T [48] 31 5.6 83.6
FAT-B3 [14] 29 4.4 83.6
SMT-S [34] 20 4.8 83.7
BiFormer-S [62] 26 4.5 83.8
SecViT-S 27 4.4 84.2
Cost Model Parmas (M) FLOPs (G) Top1-acc (%)
base model $\sim 9.0$G Swin-S [36] 50 8.7 83.0
ConvNeXt-S [37] 50 8.7 83.1
CrossFormer-B [53] 52 9.2 83.4
NAT-S [18] 51 7.8 83.7
Quadtree-B-b4 [46] 64 11.5 84.0
ScaleViT-B [58] 81 8.6 84.1
MOAT-1 [56] 42 9.1 84.2
InternImage-S [50] 50 8.0 84.2
DaViT-S [11] 50 8.8 84.2
BiFormer-B [62] 57 9.8 84.3
MViTv2-B [30] 52 10.2 84.4
CMT-B [16] 46 9.3 84.5
iFormer-B [45] 48 9.4 84.6
SE-CoTNetD-152 [31] 56 26.5 84.6
SecViT-B 55 9.7 85.0
large model $\sim 18.0$G DeiT-B [47] 86 17.5 81.8
Swin-B [36] 88 15.4 83.3
LITv2 [40] 87 13.2 83.6
CrossFormer-L [53] 92 16.1 84.0
Ortho-L [25] 88 15.4 84.2
CSwin-B [12] 78 15.0 84.2
SMT-L [34] 81 17.7 84.6
DaViT-B [11] 88 15.5 84.6
SG-Former-B [15] 78 15.6 84.7
iFormer-L [45] 87 14.0 84.8
CMT-L [16] 75 19.5 84.8
InterImage-B [50] 97 16.0 84.9
MaxViT-B [48] 120 23.4 84.9
GC-ViT-B [20] 90 14.8 85.0
SecViT-L 96 18.1 85.6
Table 1: Comparison with the state-of-the-art on ImageNet-1K classification.
Comparison with SOTA.
We compare our SecViT with numerous state-of-the-art models, the results are shown in Tab.1. SecViT consistently outperforms preceding models across all scales. Notably, SecViT-S attains a Top1-accuracy of 84.2% with a mere 15M parameters and 2.5G FLOPs. For larger models, SecViT-L achieves a Top1-accuracy of 85.6% with 96M parameters and 18.1G FLOPs.
Model Parmas (M) FLOPs (G) Throughput (imgs/s) Top1-acc (%)
Swin-T [36] 29 4.5 1723 81.3
SEC-Swin-T 29 4.9 1482 83.6(+2.3)
Swin-S [36] 50 8.8 1062 83.0
SEC-Swin-S 50 9.3 804 84.8(+1.8)
Model Parmas (M) FLOPs (G) Throughput (imgs/s) Top1-acc (%)
FasterViT-0 [19] 31 3.3 3551 82.1
SEC-FasterViT-0 25 3.6 3218 82.6(+0.5)
FasterViT-1 [19] 53 5.3 2619 83.2
SEC-FasterViT-1 41 5.6 2486 83.8(+0.6)
Table 2: Strict comparison with the baslines on ImageNet-1K classfication. The speed of the models are measured on A100 GPU with the batch size of 64.
Strict Comparison with baselines.
We select two baselines: a general-purpose backbone, Swin-Transformer [36], and a efficiency-oriented backbone, FasterViT [19] to make a comparison with our SecViT. In the comparison models (SEC-Swin and SEC-FasterViT), we merely substitute the attention mechanism in the original Swin-Transformer and FasterViT with our Semantic Equitable Clustering based Self-Attention and without introducing any other modules (such as CPE, Conv Stem, etc.). As shown in Tab. 2, the simple replacement of the attention mechanism yields significant advantages in both performance and efficiency. Specifically, SEC-Swin achieves or even surpasses a 2.0% improvement over Swin in Tiny and Small model sizes. Meanwhile, SEC-FasterViT attains higher accuracy than FasterViT while utilizing fewer parameters.
4.2 Object Detection and Instance Segmentation
Settings.
We utilize MMDetection [6] to implement Mask-RCNN [21], Cascade Mask R-CNN [3], and RetinaNet [32] to evaluate the performance of the SecViT. For Mask R-CNN and Cascade Mask R-CNN, we follow the commonly used “3×+3$\times$+3 × + MS” setting, and for Mask R-CNN and RetinaNet, we apply the “1×1$\times$1 ×” setting. Following previous works [36, 12, 62], during training, we resize the images such that the shorter side is 800 pixels while keeping the longer side within 1333 pixels.
Backbone Params (M) FLOPs (G) Mask R-CNN 3×3$\times$3 ×+MS
$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
Focal-T [57] 49 291 47.2 69.4 51.9 42.7 66.5 45.9
NAT-T [18] 48 258 47.8 69.0 52.6 42.6 66.0 45.9
GC-ViT-T [20] 48 291 47.9 70.1 52.8 43.2 67.0 46.7
MPViT-S [28] 43 268 48.4 70.5 52.6 43.9 67.6 47.5
SMT-S [34] 40 265 49.0 70.1 53.4 43.4 67.3 46.7
CSWin-T [12] 42 279 49.0 70.7 53.7 43.6 67.9 46.6
InternImage-T [50] 49 270 49.1 70.4 54.1 43.7 67.3 47.3
SecViT-S 45 259 51.3 72.4 55.9 45.4 69.7 48.8
ConvNeXt-S [37] 70 348 47.9 70.0 52.7 42.9 66.9 46.2
NAT-S [18] 70 330 48.4 69.8 53.2 43.2 66.9 46.4
Swin-S [36] 69 359 48.5 70.2 53.5 43.3 67.3 46.6
InternImage-S [50] 69 340 49.7 71.1 54.5 44.5 68.5 47.8
SMT-B [34] 52 328 49.8 71.0 54.4 44.0 68.0 47.3
CSWin-S [12] 54 342 50.0 71.3 54.7 44.5 68.4 47.7
SecViT-B 73 368 52.5 73.4 57.6 46.1 70.6 49.8
Backbone Params (M) FLOPs (G) Cascade Mask R-CNN 3×3$\times$3 ×+MS
$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
NAT-T [18] 85 737 51.4 70.0 55.9 44.5 67.6 47.9
GC-ViT-T [20] 85 770 51.6 70.4 56.1 44.6 67.8 48.3
SMT-S [34] 78 744 51.9 70.5 56.3 44.7 67.8 48.6
UniFormer-S [29] 79 747 52.1 71.1 56.6 45.2 68.3 48.9
Ortho-S [25] 81 755 52.3 71.3 56.8 45.3 68.6 49.2
HorNet-T [43] 80 728 52.4 71.6 56.8 45.6 69.1 49.6
CSWin-T [12] 80 757 52.5 71.5 57.1 45.3 68.8 48.9
SecViT-S 83 738 53.9 72.6 58.6 46.7 70.1 51.1
Swin-S [36] 107 838 51.9 70.7 56.3 45.0 68.2 48.8
NAT-S [18] 108 809 51.9 70.4 56.2 44.9 68.2 48.6
GC-ViT-S [20] 108 866 52.4 71.0 57.1 45.4 68.5 49.3
DAT-S [54] 107 857 52.7 71.7 57.2 45.5 69.1 49.3
CSWin-S [12] 92 820 53.7 72.2 58.4 46.4 69.6 50.6
UniFormer-B [29] 107 878 53.8 72.8 58.5 46.4 69.9 50.4
SecViT-B 111 847 55.1 73.9 59.8 47.6 71.5 51.6
Table 3: Comparison with other backbones using "$3\times+\mathrm{MS}$“ schedule on COCO.
Results.
Tab. 3 and Tab. 4 show the results of SecViT with different detection frameworks. The results show that SecViT performs better than its counterparts in all comparisons. For “3×+3$\times$+3 × +MS” schedule, SecViT outperforms the recent InternImage [50] for +2.2 box AP and +1.7 mask AP with the framework of Mask R-CNN. Regarding the Cascade Mask R-CNN, SecViT still has very good performance. As for “1×1$\times$1 ×” schedule, SecViT also demonstrate very impressive performance.
Backbone Params (M) FLOPs (G) Mask R-CNN 1×1$\times$1 × Params (M) FLOPs (G) RetinaNet 1×1$\times$1 ×
$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$ $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{b}_{S}$ $AP^{b}_{M}$ $AP^{b}_{L}$
PVTv2-B1 [52] 33 243 41.8 54.3 45.9 38.8 61.2 41.6 23 225 41.2 61.9 43.9 25.4 44.5 54.3
FAT-B2 [14] 33 215 45.2 67.9 49.0 41.3 64.6 44.0 23 196 44.0 65.2 47.2 27.5 47.7 58.8
SecViT-T 34 221 47.7 69.5 52.6 42.9 66.4 46.1 24 202 45.6 66.8 48.9 28.8 49.8 60.7
Swin-T [36] 48 267 43.7 66.6 47.7 39.8 63.3 42.7 38 248 41.7 63.1 44.3 27.0 45.3 54.7
CMT-S [16] 45 249 44.6 66.8 48.9 40.7 63.9 43.4 44 231 44.3 65.5 47.5 27.1 48.3 59.1
MPViT-S [28] 43 268 46.4 68.6 51.2 42.4 65.6 45.7 32 248 45.7 57.3 48.8 28.7 49.7 59.2
STViT-S [24] 44 252 47.6 70.0 52.3 43.1 66.8 46.5
FAT-B3 [14] 49 47.6 69.7 52.3 43.1 66.4 46.2 39 45.9 66.9 49.5 29.3 50.1 60.9
BiFormer-S [62] 47.8 69.8 52.3 43.2 66.8 46.5 45.9 66.9 49.4 30.2 49.6 61.7
SecViT-S 45 259 49.7 70.8 54.7 44.2 68.1 48.0 35 240 48.1 69.3 51.8 31.3 53.0 63.6
Swin-S [36] 69 359 45.7 67.9 50.4 41.1 64.9 44.2 60 339 44.5 66.1 47.4 29.8 48.5 59.1
ScalableViT-B [58] 95 349 46.8 68.7 51.5 42.5 65.8 45.9 85 330 45.8 67.3 49.2 29.9 49.5 61.0
InternImage-S [50] 69 340 47.8 69.8 52.8 43.3 67.1 46.7
CSWin-S [12] 54 342 47.9 70.1 52.6 43.2 67.1 46.2
STViT-B [24] 70 359 49.7 71.7 54.7 44.8 68.9 48.7
SecViT-B 73 368 51.2 72.7 56.4 45.4 69.7 49.0 63 349 48.9 70.2 52.7 32.0 53.6 64.6
Swin-B [36] 107 496 46.9 69.2 51.6 42.3 66.0 45.5 98 477 45.0 66.4 48.3 28.4 49.1 60.6
Focal-B [57] 110 533 47.8 70.2 52.5 43.2 67.3 46.5 101 514 46.3 68.0 49.8 31.7 50.4 60.8
MPViT-B [28] 95 503 48.2 70.0 52.9 43.5 67.1 46.8 85 482 47.0 68.4 50.8 29.4 51.3 61.5
CSwin-B [12] 97 526 48.7 70.4 53.9 43.9 67.8 47.3
InternImage-B [50] 115 501 48.8 70.9 54.0 44.0 67.8 47.4
SecViT-L 115 546 51.9 73.4 57.0 46.1 70.4 49.6 105 527 49.9 71.2 53.7 33.0 54.5 66.1
Table 4: Comparison to other backbones using "1×1$\times$1 ד schedule on COCO.
4.3 Semantic Segmentation
Settings.
We utilize Semantic FPN [27] and UperNet [55] to validate our SecViT’s performance, implementing these frameworks via MMSegmentation [8]. We follow the PVT’s [51] training settings for Semantic FPN and train the model for 80k iterations. All models use an input resolution of $512\times 512$, and during testing, the image’s shorter side is resized to 512 pixels. Different from FPN, the UperNet is trained for 160K iterations and follows Swin’s [36] settings. We employ the AdamW optimizer with a weight decay of 0.01 and a 1500 iteration warm-up.
Semantic FPN
Backbone Params(M) FLOPs(G) mIoU(%)
PVTv2-B1 [52] 18 34 42.5
FAT-B2 [14] 17 32 45.4
EdgeViT-S [39] 17 32 45.9
SecViT-T 18 34 47.2
DAT-T [54] 32 198 42.6
CSWin-T [12] 26 202 48.2
Shuted-S [44] 26 183 48.2
FAT-B3 [14] 33 179 48.9
SecViT-S 29 177 49.5
DAT-S [54] 53 320 46.1
RegionViT-B+ [5] 77 459 47.5
UniFormer-B [29] 54 350 47.7
CSWin-S [12] 39 271 49.2
SecViT-B 57 288 50.7
DAT-B [54] 92 481 47.0
CrossFormer-L [53] 95 497 48.7
CSWin-B [12] 81 464 49.9
SecViT-L 99 471 52.0
UperNet
Backbone Params(M) FLOPs(G) mIoU(%)
DAT-T [54] 60 957 45.5
NAT-T [18] 58 934 47.1
InternImage-T [50] 59 944 47.9
MPViT-S [28] 52 943 48.3
SMT-S [34] 50 935 49.2
SecViT-S 56 933 50.6
DAT-S [54] 81 1079 48.3
SMT-B [34] 62 1004 49.6
InterImage-S [50] 80 1017 50.2
MPViT-B [28] 105 1186 50.3
CSWin-S [12] 65 1027 50.4
SecViT-B 84 1046 52.0
Swin-B [36] 121 1188 48.1
GC ViT-B [20] 125 1348 49.2
DAT-B [54] 121 1212 49.4
InternImage-B [50] 128 1185 50.8
CSWin-B [12] 109 1222 51.1
SecViT-L 130 1256 53.6
Table 5: Comparison with the state-of-the-art on ADE20K.
Results.
The results of semantic segmentation can be found in the Tab. 5. All the FLOPs are measured with the input resolution of $512\times 2048$, except the group of the SecViT-T, which are measured with the input resolution of $512\times 512$. SecViT achieves the best performance in all settings. Specifically, with the framework of Sementic FPN, our SecViT-S outperforms FAT-B3 for +0.6 mIoU. And SecViT-B surpasses CSWin-S by a large margin (+1.5 mIoU). For the framework of Upernet, SecViT-S surpasses the recent SMT-S by +1.4 mIoU. SecViT-B and SecViT-L also surpass all their counterparts.
4.4 Ablation Study
Comparison with Swin-Transformer.
As mentioned earlier, based on the Swin-T architecture, we directly replaced WSA/SWSA with our SECA. As shown in Tab. 6, this resulted in a significant performance improvement (classification accuracy increased by 2.3%).
Comparison with CSWin-Transformer.
As shown in Tab. 6, based on the SecViT-T architecture, we replace SECA with CSWSA from CSwin-Transformer. This leads to a significant drop in classification performance (a decrease of 1.2%). This fully demonstrates the effectiveness of our proposed token clustering method.
LCE.
LCE is a simple depth-wise convolution module used to further enhance the model’s ability to model local information. We conduct ablation experiments on LCE and find that it contributes to a certain improvement in the model’s performance. The results are shown in Tab. 6. Specifically, it boosts the model’s classification accuracy by +0.2.
Model Params(M) FLOPs(G) Top1-acc(%)
Swin-T 29 4.5 81.3
SEC-Swin-T 29 4.9 83.6(+2.3)
SecViT-T 15 2.5 82.6
$\xrightarrow{}$ [12] 15 2.5 81.4(-1.2)
w/o LCE 15 2.5 82.4
w/o CPE 15 2.5 82.4
w/o Conv Stem 14 2.1 82.3
Table 6: Ablation Study.
CPE.
The CPE proposed in CPVT [7] is a plug-and-play, highly portable positional encoding method, widely adopted by many methods to provide positional information to the model. It consists of a residual block containing only a 3x3 depth-wise convolution. Ablation results on CPE in Tab. 6 indicate that it offers some performance gains to the model (approximately +0.2 increase in classification accuracy).
num FLOPs(G) Top1-acc(%)
64, 16, 4, – 4.3 84.0
64, 16, 2, – 4.3 84.1
64, 8, 2, – 4.3 84.1
32, 8, 2, – 4.4 84.2
Table 7: Effect of number of clusters.
Conv Stem.
The Conv Stem is utilized in the early stages of the model to extract more refined local features. As indicated by the data in Tab. 6, the Conv Stem enhances the performance of the model to some extent in classification task. Specifically, the Conv Stem improves the classification accuracy by +0.3.
Effect of number of clusters.
In Tab. 7, we validate the impact of different cluster numbers chosen at each of the first three stages of the model on performance. For SecViT-S, we select 32, 8, and 2 clusters for the first three stages, respectively. This configuration achieves a classification accuracy of 84.2%. As we gradually increase the number of clusters at each stage, the model’s overall computation decreases, with a slight drop in accuracy. This demonstrates the robustness of Semantic Clustering with respect to the number of clusters.
4.5 Visualization of Semantic Clustering
Figure 4: Visualization results for SecViT-B. The first row shows the original images, the second row displays the clustering results of the first stage (32 clusters), the second row displays the clustering results of the second stage (8 clusters), and the third row shows the clustering results of the third stage (2 clusters).
To further understand how our semantic equitable clustering works, we visualized the clustering results of some samples. As shown in Fig. 4, we selected the first, second, and third stages of the pre-trained SecViT-B to observe the clustering results of different tokens. In the figure, patches with the same color belong to the same cluster. It is evident that in the first and second stages, our model tends to capture detailed information, while in the third stage, it tends to capture semantic information.
5 Conclusion
We propose a simple and straightforward clustering method for Vision Transformers—Semantic Equitable Clustering (SEC). This method assigns each token to a cluster by calculating the similarity between each token and a global token, then sorting them accordingly. Our clustering method takes into account the semantic information contained in the tokens, completes the clustering in a single step, and ensures an equal number of tokens in each cluster, facilitating efficient parallel processing on modern GPUs. Based on Semantic Equitable Clustering, we designed SecViT, a versatile vision backbone that achieves impressive results across various vision tasks, including image classification, object detection, instance segmentation, and semantic segmentation.
Supplementary Material
Model Blocks Channels Heads Ratios Params(M) FLOPs(G)
SecViT-T [2, 2, 9, 2] [64, 128, 256, 512] [2, 4, 8, 16] 3 15 2.5
SecViT-S [3, 4, 18, 4] [64, 128, 256, 512] [2, 4, 8, 16] 3 27 4.4
SecViT-B [4, 8, 26, 8] [80, 160, 320, 512] [5, 5, 10, 16] 3 55 9.7
SecViT-L [4, 8, 26, 8] [112, 224, 448, 640] [7, 7, 14, 20] 3 96 18.1
SEC-Swin-T [2, 2, 6, 2] [96, 192, 384, 768] [3, 6, 12, 24] 4 29 4.9
SEC-Swin-S [2, 2, 18, 2] [96, 192, 384, 768] [3, 6, 12, 24] 4 50 9.3
SEC-FasterViT-0 [2, 3, 6, 5] [64, 128, 256, 512] [–, –, 8, 16] 4 25 3.6
SEC-FasterViT-1 [1, 3, 8, 5] [80, 160, 320, 640] [–, –, 8, 16] 4 41 5.6
Table 8: Detailed Architectures of our models.
Appendix A Architecture Details
SecViT’s architecture details are illustrated in Table 8. In SecViT, we adopt four $3\times 3$ convolutions to embed the input image into tokens, batch normalization and GELU are used after each convolution. $3\times 3$ convolutions with stride 2 are used between stages to reduce the feature resolution. $3\times 3$ DWConvs are adopted in CPE. While $5\times 5$ depth-wise convolutions are adopted for LCE.
For SEC-Swin, we adhere to the design principles of the Swin-Transformer [36] without using additional structures such as CPE or Conv Stem.
For SEC-FasterViT, we adhere to the design principles of FasterViT [19]. We use convolutional for fast inference speed in the first two stages of the model and use our Semantic Equitable Clustering based Self-Attention (SECA) in the latter two stages.
For all models, we set the number of clusters in the first three stages to 32, 8, and 2, respectively.
Appendix B Experimental Settings
ImageNet Image Classification.
We adopt the training strategy proposed in DeiT [47], with the only supervision is cross entropy loss. All of our models are trained from scratch for 300 epochs with the input resolution of $224\times 224$. The AdamW is used with a cosine decay learning rate scheduler and 5 epochs of linear warm-up. The batch-size is set to 1024, respectively. We apply the same data augmentation and regularization used in DeiT [47], including RandAugment [9] (randm9-mstd0.5-inc1) , Mixup [60] (prob = 0.8), CutMix [59] (prob = 1.0), Random Erasing (prob = 0.25), and Exponential Moving Average (EMA) [41]. The maximum rates of increasing stochastic depth [23] are set to 0.1/0.15/0.4/0.5 for SecViT-T/S/B/L.
COCO Object Detection and Instance Segmentation.
We apply RetinaNet [32], Mask-RCNN [21], and Cascaded Mask R-CNN [3] as the detection frameworks based on the MMDetection [6]. All of our models are trained with “1 $\times$" (12 training epochs) and “3 $\times$ +MS" (36 training epochs with multi-scale training) settings. For the “1 $\times$" setting, images are resized to the shorter side of 800 pixels while the longer side is within 1333 pixels. For the “3 $\times$ +MS", multi-scale training strategy is used to randomly resize the shorter side of images between 480 to 800 pixels. For both frameworks, we use the AdamW with the initial learning rate of 1e-4. For RetinaNet, we set the weight decay to 1e-4. While for Mask-RCNN and Cascaded Mask R-CNN, we set it to 5e-2.
ADE20K Semantic Segmentation.
Based on MMSegmentation [8], we implement UperNet [55] and SemanticFPN [27] to evaluate our models’ performance on semantic segmentation. For UperNet, we follow the previous setting of Swin-Transformer [36] and train the model for 160k iterations with the input size of $512\times 512$. For SemanticFPN, we also use the input resolution of $512\times 512$ but train the models for 80k iterations.
Appendix C Limitations and Future Work
While the proposed Semantic Equitable Clustering have demonstrated significant efficiency for clustering tokens and impressive performance on many vision tasks, there are still several limitations that need to be addressed in our future work. First, the computational constraints prevented us from experimenting with larger models and datasets such as ImageNet-21k. Exploring the potential of Semantic Equitable Clustering on large-scale datasets and larger model could provide further insights into its scalability and efficiency. In the future, we will strive to validate the performance of our SecViT on large datasets and with larger models.
Appendix D Broader Impact Statement
The development of the Semantic Equitable Clustering (SEC) has the potential to impact the field of computer vision by offering a efficient and semantic-related approach to group tokens in Vision Transformer. By clustering the tokens, SEC reduces computational load and improves the efficiency of vision models, which could lead to broader applications.
The proposed SecViT is a general vision backbone that can be applied on different vision tasks, e.g., image classification, object detection, instance segmentation, and semantic segmentation. It has no direct negative social impact. Possible malicious uses of SecViT as a general-purpose backbone are beyond the scope of our study to discuss.
Appendix E Code
We provide the code of our Semantic Equitable Clustering base Self-Attention.
1class SECA(nn.Module):
2
3 def __init__(self, embed_dim, num_heads, num_clusters):
4 super().__init__()
5 self.num_clusters = num_clusters
6 self.embed_dim = embed_dim
7 self.num_heads = num_heads
8 self.head_dim = embed_dim // num_heads
9 self.scaling = self.head_dim ** -0.5
10 self.qkv = nn.Conv2d(embed_dim, embed_dim*3, 1, bias=True)
11 self.out_proj = nn.Conv2d(embed_dim, embed_dim, 1, bias=True)
12
13 def sort_sim(self, q: torch.Tensor, k: torch.Tensor):
14 ’’
15 q: b n (h w) d
16 k: b n (h w) d
17 ’’
18 cls_token = q.mean(2, keepdim=True) # (b n 1 d)
19 cls_token = F.normalize(cls_token, dim=-1)
20 k = F.normalize(k, dim=-1)
21 cos_sim = cls_token @ k.transpose(-1, -2) # (b n 1 (h w))
22 sort_indices = torch.argsort(cos_sim, dim=-1)
23 return sort_indices.squeeze(2) # (b n (h w))
24
25 def sort_tensor(self, indices: torch.Tensor, target: torch.Tensor):
26 ’’
27 indices: (b n l)
28 target: (b n l d)
29 ’’
30 sorted_tensor = torch.gather(target, dim=-2, index=indices.unsqueeze(-1).expand_as(target))
31 return sorted_tensor
32
33 def restore_tensor(self, indices: torch.Tensor, target: torch.Tensor):
34 ’’
35 indices: (b n l)
36 target: (b n l d)
37 ’’
38 restored_tensor = torch.gather(target, dim=-2, index=indices.sort(dim=-1).indices.unsqueeze(-1).expand_as(target))
39 return restored_tensor
40
41 def forward(self, x: torch.Tensor):
42 ’’
43 x: (b c h w)
44 ’’
45 bsz, _, h, w = x.size()
46 l = h * w
47 assert l % self.num_clusters == 0
48 qkv = self.qkv(x) # (b 3*c h w)
49
50 q, k, v = rearrange(qkv, ’b (m n d) h w -> m b n (h w) d’, m=3, n=self.num_heads)
51 # q, k, v -> b n (h w) d
52 k = k * self.scaling
53
54 if self.num_clusters > 1:
55 indices = self.sort_sim(k, k) # (b n (h w))
56 q = self.sort_tensor(indices, q) # (b n (h w) d)
57 k = self.sort_tensor(indices, k) # (b n (h w) d)
58 v = self.sort_tensor(indices, v) # (b n (h w) d)
59
60 q_l = rearrange(q, ’b n (c l) d -> b n c l d’, c=self.num_clusters)
61 k_l = rearrange(k, ’b n (c l) d -> b n c l d’, c=self.num_clusters)
62 v = rearrange(v, ’b n (c l) d -> b n c l d’, c=self.num_clusters)
63 attn = torch.softmax(q_l @ k_l.transpose(-1, -2), dim=-1) # (b n c l l)
64 v = attn @ v # (b n c l d)
65
66 v = rearrange(v, ’b n c l d -> b n (c l) d’)
67 v = self.restore_tensor(indices, v) # (b n (h w) d)
68 v = rearrange(v, ’b n (h w) d -> b (n d) h w’, h=h, w=w)
69 else:
70 attn = torch.softmax(q @ k.transpose(-1, -2), dim=-1) # (b n (h w) (h w))
71 v = attn @ v # (b n (h w) d)
72 v = rearrange(v, ’b n (h w) d -> b (n d) h w’, h=h, w=w)
73 return self.out_proj(res)

### v2

@@
-Semantic Equitable Clustering: A Simple, Fast and Effective Strategy for Vision Transformer
+Semantic Equitable Clustering: A Simple and Effective Strategy for Clustering Vision Tokens
 Abstract
@@
 Capitalizing on SEC, we propose a versatile vision backbone, SECViT.
-Comprehensive experiments in image classification, object detection, instance segmentation, and semantic segmentation validate to the effectiveness of SecViT.
-Remarkably, SecViT attains an impressive 84.2% image classification accuracy with only 27M parameters and 4.4G FLOPs, without the need for for additional supervision or data.
-Code will be available at https://github.com/qhfan/SecViT.
+Comprehensive experiments in image classification, object detection, instance segmentation, and semantic segmentation validate the effectiveness of SECViT.
+Moreover, SEC can be conveniently and swiftly applied to multimodal large language models (MLLM), such as LLaVA, to serve as a vision language connector, effectively accelerating the model’s efficiency while maintaining unchanged or better performance.
 1 Introduction
@@
 The resulting clusters, containing an equal number of tokens, can be processed in parallel by the GPU efficiently.
-Figure 2: Top-1 accuracy v.s. FLOPs on ImageNet-1K of resent SOTA models. SecViT outperforms all the counterparts in all settings.
-[…]
+Figure 2: Left: Top-1 accuracy v.s. FLOPs on ImageNet-1K of recent SOTA models. Right: Comparison among different vision language connectors on LLaVA-1.5
+Building upon Semantic Equitable Clustering (SEC), we introduce the Semantic Equitable Clustering Vision Transformer (SECViT), a versatile vision backbone that is adaptable to a wide spectrum of downstream tasks.
+As demonstrated in Fig.
+2, SECViT exhibits significant performance improvements compared to previous state-of-the-art (SOTA) models.
+Impressively, SECViT attains an accuracy of 84.3% utilizing merely 4.6GFLOPS, without the need for additional training data or supervision.
+This superior performance is maintained across different model scales.
+Furthermore, SECViT proves its proficiency in downstream tasks, including but not limited to, object detection, instance segmentation, and semantic segmentation.
+Beyond vision tasks, we also apply SEC to multimodal large language models (MLLM) such as LLaVA-1.5 [39] to serve as an efficient vision language connector.
+Specifically, we use SEC to cluster the vision tokens, and then merge all the tokens at corresponding positions within each cluster into a single token.
+Experiments demonstrate that this approach significantly enhances the efficiency of LLaVA-1.5 while improving the model’s performance.
 2 Related Works
 Vision Transformer.
-The Vision Transformer (ViT) [13] is considered a powerful visual architecture due to its excellent visual modeling capabilities.
+The Vision Transformer (ViT) [13] is considered a powerful visual architecture.
 Many works have improved the Vision Transformer, including enhancing its training efficiency and reducing its computational cost [41, 12, 46, 27, 66].
-DeiT uses distillation loss and incorporates extensive data augmentation methods into the ViT training process.
-This increases the model’s data utilization efficiency, allowing it to train an effective ViT with less data.
-Hierarchical structures represented by PVT [51, 52, 45, 16, 54] reduce the number of tokens in global attention by downsampling the keys and values (KV), thereby keeping the computational cost of global attention within an acceptable range.
-In addition to using downsampling to reduce the number of tokens, some methods directly prune tokens based on their importance, retaining important tokens while discarding or merging unimportant ones [15, 42].
+DeiT [46] uses distillation loss and incorporates extensive data augmentation methods into the ViT training process.
+Hierarchical structures represented by PVT [48, 49, 44, 18, 54] reduce the number of tokens in global attention by downsampling the keys and values (KV), thereby low the computational cost.
+In addition to them, some methods directly prune tokens based on their importance, retaining important tokens [43, 35].
 This reduces the number of tokens and subsequently lowers the computational cost of the model.
@@
 This significantly reduces the model’s computational cost.
-In contrast, CSWin [12] divides tokens into a crisscrossing structure, which also reduces the model’s computational cost.
 In addition to dividing tokens into small windows along the spatial dimension, DaViT [11] also splits channels into multiple groups along the channel dimension.
 Unlike the above methods that only consider positional information for grouping, DGT [40] takes semantic information into account by using k-means clustering to group the queries.
+Vision Language Connector.
+The vision language connector is a critical component in MLLMs [39, 3, 26].
+It aligns vision tokens with language tokens.
+Typical vision language connectors include MLP [39], Resampler [1], C-Abstractor [3], and others.
+Although MLP performs well, it introduces a significant number of vision tokens, which hampers the model’s efficiency.
+On the other hand, connectors like Resampler improve the model’s efficiency, but at the cost of reduced performance.
+Unlike these methods, our proposed SEC consider the semantic information of each token and significantly enhances the model’s efficiency while maintaining its performance.
+Figure 3: (a) Illustration of SECViT (b) Applying SEC to vision language connector. (c) Illustration of Semantic Equitable Clustering for ViT and vision language connector.
 3 Method
@@
 3(a).
-For each input image $x\in\mathbb{R}^{3\times H\times W}$, we first process it by a patch embedding to downsample it.
-Then we obtain the tokens with a shape of $C_{1}\times\frac{H}{4}\times\frac{W}{4}$.
-Similar to previous work [36, 51, 12], we adopt a hierarchical structure to process tokens.
-Our model consists of four stages with downsampling factors of $\frac{1}{4}$, $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$, respectively.
+SECViT consists of four stages with downsampling factors of $\frac{1}{4}$, $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$, respectively.
 This structural design facilitates downstream tasks, such as object detection, in constructing feature pyramids.
-A SecViT block is composed of three modules: Conditional Positional Encoding (CPE) [7], Semantic Equitable Clustering based Self-Attention (SECA), and classical Feed-Forward Network (FFN) [49].
-A complete SecViT block can be represented by Eq.
-1:
-$\centering\begin{split}&X={\rm CPE(}X_{in}{\rm)}+X_{in},\\ &Y={\rm SECA(}{\rm LN(}X{\rm))}+X,\\ &Z={\rm FFN(}{\rm LN(}Y{\rm))}+Y.\end{split}\@add@centering$ (1)
+A SECViT block is composed of three modules.
 For each block, the input tensor $X_{in}\in\mathbb{R}^{C\times H\times W}$ is fed into the CPE to introduce the positional information.
@@
 The final FFN is utilized to integrate channel-wise information of tokens.
+Beyond the design of the backbone, we also utilize SEC in the design of the vision language connector in MLLM [39].
+For the vision tokens output by ViT, we use SEC to cluster the tokens.
+For each position corresponding to a cluster, we use attentive pooling to merge them into a single token, thereby reducing the number of vision tokens.
+The process is shown in Fig.
+3(b).
 3.2 Semantic Equitable Clustering
@@
 3(c).
-Figure 3: (a) The overall architecture of our Semantic Equitable Clustering Vision Transformer.
-(b) The composition of each SecViT block.
-(c) Illustration of our Semantic Equitable Clustering.
-SEC determines the clustering rules for tokens by calculating the cosine similarity between all tokens and the global token.
 Single Clustering Center Related to Semantics.
@@
 This is because, in most vision foundation models, the output of the average pool is assumed to contain the richest semantic information and is thus used for classification [41, 12, 6, 14].
-ToMe [2] indicates that the $K$ contains the richest information from each token.
-Therefore, we choose the $K$ as the basis for clustering.
 Specifically, the process for determining the cluster center is shown in Eq.
-2:
-$\centering\begin{split}Q=W_{Q}X,&K=W_{K}X,V=W_{V}X,\\ &k_{c}={\rm Avg}(K).\end{split}\@add@centering$ (2)
+1:
+$\centering\begin{split}Q=W_{Q}X,&K=W_{K}X,V=W_{V}X,\\ &k_{c}={\rm Pool}(K).\end{split}\@add@centering$ (1)
 Where $W_{K}$ is a learnable matrix. $k_{c}$ is the determined cluster center. $X$ is the set of input tokens.
@@
 The specific process is shown in Eq.
-3:
+2:
 $\centering\begin{split}&sim=\frac{K\cdot k_{c}}{||K||\cdot||k_{c}||},\\ &idx={\rm argsort}(sim),\\ Q^{*}=Q[idx]&,K^{*}=K[idx],V^{*}=V[idx].\end{split}\@add@centering$ (2)
@@
 The obtained $Q^{*}$, $K^{*}$, and $V^{*}$ from the previous step have been sorted based on their distances to the cluster center.
-At this point, we directly group them, so tokens with similar distances to the cluster center are classified into the same cluster.
+For the design of vision backbone, we directly group them, so tokens with similar distances to the cluster center are classified into the same cluster.
 This allows us to directly control an equal number of tokens in each cluster.
@@
 3(c) and denoted as follows:
-$\centering\begin{split}Q_{c}=Q_{c\times N_{c}:\left(c+1\right)N_{c}}^{*},K_{c}=K_{c\times N_{c}:\left(c+1\right)N_{c}}^{*},K_{c}=K_{c\times N_{c}:\left(c+1\right)N_{c}}^{*}.\end{split}\@add@centering$ (4)
-where $N_{c}$ is the basic token number of each cluster for equal partition and $c$ is the index of the cluster
+$\centering\begin{split}Q_{m}=Q^{*}[{m\times N:\left(m+1\right)N}],\\ K_{m}=K^{*}[{m\times N:\left(m+1\right)N}],\\ V_{m}=V^{*}[{m\times N:\left(m+1\right)N}].\end{split}\@add@centering$ (3)
+where $N$ is the basic token number of each cluster for equal partition and $m$ is the index of the cluster
 Based on the above steps, we have completed the clustering process that captures semantic information in the image with minimal sorting cost.
 Moreover, compared to k-means, we have achieved equi-partitioning of each cluster.
-After clustering is completed, we apply standard Self-Attention to the tokens within each cluster (SECA), thereby completing the interaction of information between tokens:
-$\centering Y_{c}=\text{Attn}\left(Q_{c},K_{c},V_{c}\right).\@add@centering$ (5)
-After the completion of SECA, to further enhance the model’s ability to capture local information, we employ a local context enhancement module to model local information, where LCE [62, 12] is a simple depth-wise convolution:
-$\centering\begin{split}X={\rm SECA}(X)+{\rm LCE}(W_{V}X),\end{split}\@add@centering$ (6)
+After clustering is completed, we apply standard Self-Attention to the tokens within each cluster, thereby completing the interaction of information between tokens:
+$\centering Y_{m}=\text{Attn}\left(Q_{m},K_{m},V_{m}\right).\@add@centering$ (4)
+For the design of vision language connector, we group the tokens according to their similarity, and the tokens within each group are interleaved, as shown in Eq. 5:
+$\centering\begin{split}Q_{n}=Q^{*}[n:N:L],\\ K_{n}=K^{*}[n:N:L],\\ V_{n}=V^{*}[n:N:L].\end{split}\@add@centering$ (5)
+in which $L$ is the token’s sequence length, $n$ is the index of group tokens.
+$N$ is the basic token number of each cluster.
+After obtaining the token groups, we perform pooling on $Q$ to effectively reduce the number of tokens input to the LLM, with each group’s output becoming a single token, as shown in Eq 6.
+$Y_{n}={\rm Attn}({\rm Pool}(Q_{n}),K_{n},V_{n}).$ (6)
+3.3 Difference between SEC and EViT.
+We use the most representative example, EViT [35], to illustrate the differences between SEC and other methods based on the similarity between the global token and other tokens.
+Pruning v.s.
+Clustering.
+Most similarity-based methods, such as EViT, are pruning methods, where tokens with low similarity to the [cls] token are merged during the forward process, thereby reducing the number of tokens and decreasing computational cost.
+In contrast, our proposed SECViT employs a clustering-based approach, performing attention operations within each cluster.
+The role of the [cls] token.
+In methods like EViT, the [cls] token serves as a measure of importance of a token.
+Each token computes its similarity to the [cls] token, with higher similarity tokens deemed more important.
+The less important tokens are abandoned.
+In contrast, in SEC, the [cls] token (obtained by average pooling over all tokens) measures similarity between tokens.
+Each token computes its similarity score to the [cls] token; tokens with similar scores are considered to be more similar and grouped into one cluster.
+Attention is calculated only within the same cluster.
+Models Params (M) FLOPs (G) Throughput (imgs/s) Acc $AP^{b}$ $AP^{m}$ mIoU
+DeiT-S 22 4.6 3204 79.8 44.5 40.1 43.0
+EViT-DeiT-S (keeprate=0.9) 22 4.0 3428 79.8 not suit not suit not suit
+SEC-DeiT-S (num_cluster=2) 22 4.3 3226 80.6 (+0.8) 47.9 (+3.4) 43.1 (+3.0) 48.0 (+5.0)
+SEC-DeiT-S (num_cluster=4) 22 4.1 3412 80.5 (+0.7) 47.7 (+3.2) 42.7 (+2.6) 47.5 (+4.5)
+SEC-DeiT-S (num_cluster=8) 22 3.9 3528 80.1 (+0.3) 46.7 (+2.2) 42.0 (+1.9) 46.4 (+3.4)
+Table 1: Comparison of EViT [35] and SEC. Inference speed are measured on the A100 GPU.
+Different adaptability to vision models/MLLMs.
+In pruning methods like EViT, during the model’s forward pass, the number of tokens gradually decreases.
+Although reducing tokens does not impact classification tasks, it prevents the feature map from being restored to its original shape.
+This makes it difficult for EViT to be directly used with classic frameworks like SemanticFPN for downstream dense prediction tasks .
+In SEC, we simply group the tokens without changing their quantity, thereby preserving the integrity of the feature map.
+This ensures that SEC can be easily applied to downstream tasks.
+In Tab.
+1, we use DeiT [46] as the baseline.
+Without introducing any other tricks, just by applying SEC to DeiT, we form SEC-DeiT and compare it with EViT-DeiT.
+We conducted image classification, object detection, instance segmentation and semantic segmentation based on SEC-DeiT-S.
+SEC not only accelerates the model but also enhances its performance.
+Beyond the vision tasks, we also compare the EViT with our SEC on MLLM, details can be found in the Tab.
+8.
 4 Experiments
-We conducted experiments on a wide range of vision tasks, including image classification on ImageNet-1k [10], object detection and instance segmentation on COCO [33], and semantic segmentation on ADE20K [61].
-All models can be trained with 8 A100 80G GPUs.
-4.1 ImageNet Classification
-Settings.
-We train our models from scratch on ImageNet-1k [10].
-For a fair comparison, we follow the training strategy in DeiT [47], with classification loss serving as the only supervision.
-The maximum rates for increasing stochastic depth [23] are set to 0.1, 0.15, 0.4, and 0.5 for SecViT-T, SecViT-S, SecViT-B, and SecViT-L, respectively.
+We first make strict comparison with hierarchical/plain baselines.
+Then we conduct experiments on a wide range of vision tasks for SECViT, including image classification, object detection, instance segmentation, and semantic segmentation.
+We also verify the role of SEC in MLLM based on LLaVA-1.5 [39].
+More details, experiments, and comparison of models’ efficiency can be found in the Appendix.
+4.1 SEC for vision models
+Model Params (M) FLOPs (G) Throughput (imgs/s) Acc $AP^{b}$ $AP^{m}$ mIoU
+DeiT-S [46] 22 4.6 3204 79.8 44.5 40.1 43.0
+EViT-DeiT-S (keeprate=0.9) 22 4.0 3428 79.8 not suit not suit not suit
+SEC-DeiT-S (num_cluster=4) 22 4.1 3412 80.5 (+0.7) 47.7 (+3.2) 42.7 (+2.6) 47.5 (+4.5)
+DeiT-B 86 17.6 1502 81.8
+SEC-DeiT-B 86 14.8 1682 82.4 (+0.6)
+Swin-T 29 4.5 1723 81.3 43.7 39.8 44.5
+SEC-Swin-T 29 4.8 1482 83.8 (+2.5) 48.3 (+4.6) 43.4 (+3.6) 49.3 (+4.8)
+Swin-S 50 8.8 1006 83.0 45.7 41.1 47.6
+SEC-Swin-S 50 9.2 804 85.0 (+2.0) 50.2 (+4.5) 44.7 (+3.6) 51.3 (+3.7)
+Table 2: Comparison with Hierarchy/Plain baselines. Inference speed are measured on the A100 GPU.
+Model Params (M) FLOPs (G) Method Pretrain epoch Acc(%)
+Swin-B [41] 88 15.4 Supervised 83.5
+ConvNeXt V2-B [53] 88 15.4 Supervised 84.3
+SEC-Swin-B 88 16.2 Supervised 85.3
+Swin-B [41] 88 15.4 SimMIM [56] 800 84.0(+0.5)
+ConvNeXt V2-B [53] 88 15.4 FCMAE [53] 800 84.6(+0.3)
+SEC-Swin-B 88 16.2 SimMIM [56] 800 85.9(+0.6)
+Table 3: Comparison with baselines on self-supervised setting.
 Cost Model Parmas (M) FLOPs (G) Top1-acc (%)
 tiny model $\sim 2.5$G PVTv2-b1 [49] 13 2.1 78.7
+TCFormer-light [63] 14 3.8 79.4
 QuadTree-B-b1 [45] 14 2.3 80.0
-RegionViT-T [5] 14 2.4 80.4
 MPViT-XS [31] 11 2.9 80.9
-VAN-B1 [17] 14 2.5 81.1
 BiFormer-T [66] 13 2.2 81.4
-Conv2Former-N [22] 15 2.2 81.5
 CrossFormer-T [50] 28 2.9 81.5
-NAT-M [18] 20 2.7 81.8
 FAT-B2 [14] 14 2.0 81.9
-QnA-T [1] 16 2.5 82.0
 GC-ViT-XT [22] 20 2.6 82.0
 SMT-T [37] 12 2.4 82.2
-SecViT-T 15 2.5 82.6
-small model $\sim 4.5$G Swin-T [36] 29 4.5 81.3
-CrossViT-15 [4] 27 5.8 81.5
-RVT-S [38] 23 4.7 81.9
+RMT-T [15] 14 2.5 82.4
+SECViT-T 15 2.5 82.7
+small model $\sim 4.5$G PS-ViT-B14 [61] 21 5.4 81.7
+DVT-T2T-ViT-19 [52] 39 6.2 81.9
 ConvNeXt-T [42] 29 4.5 82.1
-Focal-T [57] 29 4.9 82.2
-MPViT-S [28] 23 4.7 83.0
+TCFormer [63] 26 5.8 82.3
 SG-Former-S [17] 23 4.8 83.2
-Ortho-S [25] 24 4.5 83.4
+StructViT-S-8-1 [28] 24 5.4 83.3
 InternImage-T [51] 30 5.0 83.5
-GC-ViT-T [20] 28 4.7 83.5
-CMT-S [16] 25 4.0 83.5
+MLLA-T [20] 25 4.2 83.5
 MaxViT-T [47] 31 5.6 83.6
@@
 BiFormer-S [66] 26 4.5 83.8
-SecViT-S 27 4.4 84.2
-Cost Model Parmas (M) FLOPs (G) Top1-acc (%)
-base model $\sim 9.0$G Swin-S [36] 50 8.7 83.0
-ConvNeXt-S [37] 50 8.7 83.1
-CrossFormer-B [53] 52 9.2 83.4
+SECViT-S 27 4.6 84.3
+base model $\sim 9.0$G ConvNeXt-S [42] 50 8.7 83.1
 NAT-S [21] 51 7.8 83.7
 Quadtree-B-b4 [45] 64 11.5 84.0
-ScaleViT-B [58] 81 8.6 84.1
 MOAT-1 [57] 42 9.1 84.2
 InternImage-S [51] 50 8.0 84.2
-DaViT-S [11] 50 8.8 84.2
+GC-ViT-S [22] 51 8.5 84.3
 BiFormer-B [66] 57 9.8 84.3
-MViTv2-B [30] 52 10.2 84.4
-CMT-B [16] 46 9.3 84.5
 iFormer-B [44] 48 9.4 84.6
-SE-CoTNetD-152 [31] 56 26.5 84.6
-SecViT-B 55 9.7 85.0
-large model $\sim 18.0$G DeiT-B [47] 86 17.5 81.8
-Swin-B [36] 88 15.4 83.3
-LITv2 [40] 87 13.2 83.6
-CrossFormer-L [53] 92 16.1 84.0
-Ortho-L [25] 88 15.4 84.2
-CSwin-B [12] 78 15.0 84.2
+FAT-B4 [14] 52 9.3 84.8
+SECViT-B 57 9.8 85.2
+large model $\sim 18.0$G CrossFormer-L [50] 92 16.1 84.0
 SMT-L [37] 81 17.7 84.6
@@
 iFormer-L [44] 87 14.0 84.8
-CMT-L [16] 75 19.5 84.8
 InterImage-B [51] 97 16.0 84.9
-MaxViT-B [48] 120 23.4 84.9
 GC-ViT-B [22] 90 14.8 85.0
-SecViT-L 96 18.1 85.6
+RMT-L [15] 95 18.2 85.5
+SECViT-L 101 18.2 85.7
+XL model $\sim 35.0$G ConvNeXt-L [42] 198 34.4 84.3
+CoAtNet-3 [10] 168 34.7 84.5
+MaxViT-L [47] 212 43.9 85.1
+GC ViT-L [22] 201 32.6 85.7
+SECViT-XL 205 36.4 86.3
 Table 4: Comparison with the state-of-the-art on ImageNet-1K classification.
-Comparison with SOTA.
-We compare our SecViT with numerous state-of-the-art models, the results are shown in Tab.1.
-SecViT consistently outperforms preceding models across all scales.
-Notably, SecViT-S attains a Top1-accuracy of 84.2% with a mere 15M parameters and 2.5G FLOPs.
-For larger models, SecViT-L achieves a Top1-accuracy of 85.6% with 96M parameters and 18.1G FLOPs.
-Model Parmas (M) FLOPs (G) Throughput (imgs/s) Top1-acc (%)
-Swin-T [36] 29 4.5 1723 81.3
-SEC-Swin-T 29 4.9 1482 83.6(+2.3)
-Swin-S [36] 50 8.8 1062 83.0
-SEC-Swin-S 50 9.3 804 84.8(+1.8)
-Model Parmas (M) FLOPs (G) Throughput (imgs/s) Top1-acc (%)
-FasterViT-0 [19] 31 3.3 3551 82.1
-SEC-FasterViT-0 25 3.6 3218 82.6(+0.5)
-FasterViT-1 [19] 53 5.3 2619 83.2
-SEC-FasterViT-1 41 5.6 2486 83.8(+0.6)
-Table 2: Strict comparison with the baslines on ImageNet-1K classfication. The speed of the models are measured on A100 GPU with the batch size of 64.
-Strict Comparison with baselines.
-We select two baselines: a general-purpose backbone, Swin-Transformer [36], and a efficiency-oriented backbone, FasterViT [19] to make a comparison with our SecViT.
-In the comparison models (SEC-Swin and SEC-FasterViT), we merely substitute the attention mechanism in the original Swin-Transformer and FasterViT with our Semantic Equitable Clustering based Self-Attention and without introducing any other modules (such as CPE, Conv Stem, etc.).
-As shown in Tab.
-2, the simple replacement of the attention mechanism yields significant advantages in both performance and efficiency.
-Specifically, SEC-Swin achieves or even surpasses a 2.0% improvement over Swin in Tiny and Small model sizes.
-Meanwhile, SEC-FasterViT attains higher accuracy than FasterViT while utilizing fewer parameters.
-4.2 Object Detection and Instance Segmentation
-Settings.
-We utilize MMDetection [6] to implement Mask-RCNN [21], Cascade Mask R-CNN [3], and RetinaNet [32] to evaluate the performance of the SecViT.
-For Mask R-CNN and Cascade Mask R-CNN, we follow the commonly used “3×+3$\times$+3 × + MS” setting, and for Mask R-CNN and RetinaNet, we apply the “1×1$\times$1 ×” setting.
-Following previous works [36, 12, 62], during training, we resize the images such that the shorter side is 800 pixels while keeping the longer side within 1333 pixels.
-Backbone Params (M) FLOPs (G) Mask R-CNN 3×3$\times$3 ×+MS
-$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
-Focal-T [57] 49 291 47.2 69.4 51.9 42.7 66.5 45.9
-NAT-T [18] 48 258 47.8 69.0 52.6 42.6 66.0 45.9
+Backbone Params (M) FLOPs (G) Mask R-CNN 1×1$\times$1 × Params (M) FLOPs (G) RetinaNet 1×1$\times$1 ×
+$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$ $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{b}_{S}$ $AP^{b}_{M}$ $AP^{b}_{L}$
+PVTv2-B1 [49] 33 243 41.8 54.3 45.9 38.8 61.2 41.6 23 225 41.2 61.9 43.9 25.4 44.5 54.3
+FAT-B2 [14] 33 215 45.2 67.9 49.0 41.3 64.6 44.0 23 196 44.0 65.2 47.2 27.5 47.7 58.8
+RMT-T [15] 33 218 47.1 68.8 51.7 42.6 65.8 45.9 23 199 45.1 66.2 48.1 28.8 48.9 61.1
+SECViT-T 34 221 47.8 69.5 52.5 43.0 66.7 46.3 24 202 45.8 66.8 49.2 29.1 49.8 60.9
+MPViT-S [31] 43 268 46.4 68.6 51.2 42.4 65.6 45.7 32 248 45.7 57.3 48.8 28.7 49.7 59.2
+MLLA-T [20] 44 255 46.8 69.5 51.5 42.1 66.4 45.0
+STViT-S [25] 44 252 47.6 70.0 52.3 43.1 66.8 46.5
+RMT-S [15] 46 262 49.0 70.8 53.9 43.9 67.8 47.4 36 244 47.8 69.1 51.8 32.1 51.8 63.5
+SECViT-S 45 262 49.9 70.9 54.7 44.6 68.3 47.7 35 240 48.4 69.4 52.0 31.3 53.3 63.8
+ScalableViT-B [59] 95 349 46.8 68.7 51.5 42.5 65.8 45.9 85 330 45.8 67.3 49.2 29.9 49.5 61.0
+InternImage-S [51] 69 340 47.8 69.8 52.8 43.3 67.1 46.7
+MLLA-S [20] 63 319 49.2 71.5 53.9 44.2 68.5 47.2
+STViT-B [25] 70 359 49.7 71.7 54.7 44.8 68.9 48.7
+SECViT-B 76 371 51.5 72.9 56.7 45.4 69.9 48.7 63 349 49.3 70.3 52.9 32.0 53.8 64.8
+Focal-B [58] 110 533 47.8 70.2 52.5 43.2 67.3 46.5 101 514 46.3 68.0 49.8 31.7 50.4 60.8
+CSwin-B [12] 97 526 48.7 70.4 53.9 43.9 67.8 47.3
+InternImage-B [51] 115 501 48.8 70.9 54.0 44.0 67.8 47.4
+MLLA-B [20] 115 502 50.5 72.0 55.4 45.0 69.3 48.6
+SECViT-L 119 550 52.0 73.5 57.3 46.3 70.6 49.8 105 527 50.2 71.4 53.9 33.2 54.5 66.3
+Table 5: Comparison to other backbones using ”1×1$\times$1 ד schedule on COCO.
+Backbone Params (M) FLOPs (G) $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
+Mask R-CNN 3×3$\times$3 ×+MS
 GC-ViT-T [22] 48 291 47.9 70.1 52.8 43.2 67.0 46.7
-MPViT-S [28] 43 268 48.4 70.5 52.6 43.9 67.6 47.5
+MLLA-T [20] 44 255 48.8 71.0 53.6 43.8 68.0 46.8
 SMT-S [37] 40 265 49.0 70.1 53.4 43.4 67.3 46.7
-CSWin-T [12] 42 279 49.0 70.7 53.7 43.6 67.9 46.6
 InternImage-T [51] 49 270 49.1 70.4 54.1 43.7 67.3 47.3
-SecViT-S 45 259 51.3 72.4 55.9 45.4 69.7 48.8
-ConvNeXt-S [37] 70 348 47.9 70.0 52.7 42.9 66.9 46.2
+RMT-S [15] 46 262 50.7 71.9 55.6 44.9 69.1 48.4
+SECViT-S 45 262 51.6 72.5 55.9 45.6 69.9 48.8
 NAT-S [21] 70 330 48.4 69.8 53.2 43.2 66.9 46.4
-Swin-S [36] 69 359 48.5 70.2 53.5 43.3 67.3 46.6
 InternImage-S [51] 69 340 49.7 71.1 54.5 44.5 68.5 47.8
 SMT-B [37] 52 328 49.8 71.0 54.4 44.0 68.0 47.3
-CSWin-S [12] 54 342 50.0 71.3 54.7 44.5 68.4 47.7
-SecViT-B 73 368 52.5 73.4 57.6 46.1 70.6 49.8
-Backbone Params (M) FLOPs (G) Cascade Mask R-CNN 3×3$\times$3 ×+MS
-$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
+MLLA-S [20] 63 319 50.5 71.8 55.2 44.9 69.1 48.2
+RMT-B [15] 73 373 52.2 72.9 57.0 46.1 70.4 49.9
+SECViT-B 75 371 52.8 73.6 57.7 46.4 70.8 49.9
+Cascade Mask R-CNN 3×3$\times$3 ×+MS
 NAT-T [21] 85 737 51.4 70.0 55.9 44.5 67.6 47.9
@@
 UniFormer-S [33] 79 747 52.1 71.1 56.6 45.2 68.3 48.9
-Ortho-S [25] 81 755 52.3 71.3 56.8 45.3 68.6 49.2
-HorNet-T [43] 80 728 52.4 71.6 56.8 45.6 69.1 49.6
-CSWin-T [12] 80 757 52.5 71.5 57.1 45.3 68.8 48.9
-SecViT-S 83 738 53.9 72.6 58.6 46.7 70.1 51.1
-Swin-S [36] 107 838 51.9 70.7 56.3 45.0 68.2 48.8
+RMT-S [15] 83 741 53.2 72.0 57.8 46.1 69.8 49.8
+SECViT-S 83 741 54.1 72.8 58.6 47.0 70.3 51.0
 NAT-S [21] 108 809 51.9 70.4 56.2 44.9 68.2 48.6
 GC-ViT-S [22] 108 866 52.4 71.0 57.1 45.4 68.5 49.3
-DAT-S [54] 107 857 52.7 71.7 57.2 45.5 69.1 49.3
 CSWin-S [12] 92 820 53.7 72.2 58.4 46.4 69.6 50.6
 UniFormer-B [33] 107 878 53.8 72.8 58.5 46.4 69.9 50.4
-SecViT-B 111 847 55.1 73.9 59.8 47.6 71.5 51.6
-Table 3: Comparison with other backbones using "$3\times+\mathrm{MS}$“ schedule on COCO.
-Results.
+RMT-B [15] 111 852 54.5 72.8 59.0 47.2 70.5 51.4
+SECViT-B 114 849 55.4 74.1 59.9 47.8 71.7 51.7
+Table 6: Comparison with other backbones using ”$3\times+\mathrm{MS}$“ schedule on COCO.
+Model Semantic FPN 80K Upernet 160K
+Params (M) FLOPs (G) mIoU (%) Params (M) FLOPs (G) mIoUss (%)
+PVTv2-B1 [49] 18 136 42.5
+VAN-B1 [19] 18 140 42.9
+RMT-T [15] 17 136 46.4
+SECViT-T 18 136 47.2 44 894 48.8
+StructViT-S [28] 26 271 46.9
+MogaNet-S [34] 29 189 47.7 55 946 49.2
+SMT-S [37] 50 935 49.2
+SGFormer-S [17] 25 205 49.0 52.5 989 49.9
+RMT-S [15] 30 180 49.4 56 937 49.8
+SECViT-S 30 180 49.6 56 936 50.6
+MogaNet-B [34] 74 1050 50.1
+InterImage-S [51] 80 1017 50.2
+StructViT-B [28] 54 529 48.5
+RMT-B [15] 57 294 50.4 83 1051 52.0
+SECViT-B 60 291 50.7 86 1048 52.2
+MogaNet-L [34] 113 1176 50.9
+CSWin-B [12] 81 464 49.9 109 1222 51.1
+MLLA-B [20] 128 1183 51.9
+SGFormer-B [17] 81 475 50.6 109 1304 52.0
+RMT-L [15] 98 482 51.4 125 1241 52.8
+SECViT-L 103 475 52.2 131 1256 53.8
+Table 7: Comparison with the state-of-the-art on ADE20K.
+Model Connector V-T Num Time Speed TextVQA GQA VQAv2 POPE MME
+LLaVA-1.5 MLP 576+1 194s 1.0×1.0$\times$1.0 × 58.2 62.0 78.5 86.1 1510.7
+LLaVA-1.5+Resampler Resampler 288+1 126s 1.5×1.5$\times$1.5 × 52.1 56.8 76.0 83.1 1393.2
+LLaVA-1.5+EViT MLP+EViT 288+1 126s 1.5×1.5$\times$1.5 × 54.6 60.0 77.9 84.3 1483.2
+LLaVA-1.5+SEC MLP+SEC 288+1 126s 1.5×1.5$\times$1.5 × 60.1 63.5 78.9 87.7 1510.7
+LLaVA-1.5+Resampler Resampler 256+1 116s 1.7×1.7$\times$1.7 × 51.6 56.0 75.2 82.7 1387.2
+LLaVA-1.5+Pool MLP+Pool 256+1 116s 1.7×1.7$\times$1.7 × 52.4 57.6 76.4 83.3 1415.5
+LLaVA-1.5+EViT MLP+EViT 256+1 116s 1.7×1.7$\times$1.7 × 52.8 59.6 77.1 83.7 1443.7
+LLaVA-1.5+SEC MLP+SEC 256+1 116s 1.7×1.7$\times$1.7 × 59.6 63.2 78.6 87.1 1505.2
+LLaVA-1.5+Resampler Resampler 192+1 102s 1.9×1.9$\times$1.9 × 50.1 55.2 74.3 82.7 1337.6
+LLaVA-1.5+EViT MLP+EViT 192+1 102s 1.9×1.9$\times$1.9 × 51.6 58.6 76.3 83.1 1427.6
+LLaVA-1.5+SEC MLP+SEC 192+1 102s 1.9×1.9$\times$1.9 × 57.7 62.7 78.4 86.7 1500.1
+LLaVA-1.5+Resampler Resampler 144+1 94s 2.1×2.1$\times$2.1 × 47.6 54.6 72.0 81.9 1293.7
+LLaVA-1.5+Pool MLP+Pool 144+1 94s 2.1×2.1$\times$2.1 × 50.0 56.2 73.6 81.9 1310.7
+LLaVA-1.5+EViT MLP+EViT 144+1 94s 2.1×2.1$\times$2.1 × 51.2 58.0 76.0 83.1 1393.6
+LLaVA-1.5+SEC MLP+SEC 144+1 94s 2.1×2.1$\times$2.1 × 56.8 62.0 78.0 86.1 1487.1
+Table 8: Comparison of different vision language connectors on LLaVA-1.5.
+“V-T Num” denotes the quantity of visual tokens.
+The computation expense is impacted by V-T Num, with larger values resulting in higher costs.
+“Speed” refers to the comparative inference velocity relative to LLaVA-1.5.
+“Time” is the average inference time.
+Inference speed are measured on the A100.
+Model LLM Connector V-T Num Res TextVQA GQA VQAv2 VisWiz SQAimg Speed ($\uparrow$)
+7B LLM
+Shikra [5] Vicuna-7B MLP 257 224 - - 77.4 - - -
+IDEFICS-9B [30] LLaMA-7B Cross Attn 257 224 - 38.4 50.9 35.5 - -
+Qwen-VL [1] Qwen-7B Resampler 256 448 - 59.3 78.8 35.2 67.1 -
+Qwen-VL-Chat [1] Qwen-7B Resampler 256 448 - 57.5 78.2 38.9 68.2 -
+LLaVA-1.5 [38] Vicuna-7B MLP 577 336 58.2 62.0 78.5 50.0 66.8 1.0×1.0$\times$1.0 ×
+LLaVA-1.5+SEC (ours) Vicuna-7B MLP+SEC 257 336 59.6 63.2 78.9 52.8 69.6 $\mathbf{1.7\times}$
+13B LLM
+InstructBLIP [9] Vicuna-13B Q-Former 32 224 - 49.5 - 33.4 63.1 -
+BLIP-2 [32] Vicuna-13B Q-Former 32 224 - 41.0 41.0 19.5 61.0 -
+LLaVA-1.5 [38] Vicuna-13B MLP 577 336 61.2 63.3 80.0 53.6 71.6 1.0×1.0$\times$1.0 ×
+LLaVA1.5+SEC (ours) Vicuna-13B MLP+SEC 257 336 62.3 64.3 80.0 54.7 72.0 $\mathbf{1.8\times}$
+Table 9: Results on General VQA tasks.
+Model LLM Connector V-T Num Res POPE MMB MM-Vet Speed ($\uparrow$)
+7B LLM
+MiniGPT-4 [65] Vicuna-7B Resampler 32 224 72.2 24.3 22.1 -
+mPLUG-Owl2 [60] LLaMA2-7B Resampler 32 224 - 49.4 - -
+LLaMA-AdapterV2 [16] LLaMA2-7B LLaMA-Adapter 257 224 - 41.0 31.4 -
+Shikra [5] Vicuna-7B MLP 257 224 - 58.8 - -
+Qwen-VL [1] Qwen-7B Resampler 256 448 - 38.2 - -
+Qwen-VL-Chat [1] Qwen-7B Resampler 256 448 - 60.6 - -
+LLaVA-1.5 [38] Vicuna-7B MLP 577 336 86.1 64.3 31.1 1.0×1.0$\times$1.0 ×
+LLaVA1.5+SEC (ours) Vicuna-7B MLP+SEC 145 336 86.1 68.4 31.7 $\mathbf{2.1\times}$
+13B LLM
+MiniGPT-4 [65] Vicuna-13B Resampler 32 224 - - 24.4 -
+BLIP-2 [32] Vicuna-13B Q-Former 32 224 85.3 - 22.4 -
+LLaVA-1.5 [38] Vicuna-13B MLP 577 336 86.2 67.7 36.1 1.0×1.0$\times$1.0 ×
+LLaVA-1.5+SEC (ours) Vicuna-13B MLP+SEC 145 336 86.4 69.2 37.3 $\mathbf{2.2\times}$
+Table 10: Results on benchmark designed for MLLMs.
+Strict Comparison with Baselines.
+We select two baselines: hierarchical backbone Swin-Transformer [41] and plain backbone DeiT [46] to make a comparison with our SEC based model.
+In the comparison models (SEC-Swin and SEC-DeiT), we merely substitute the attention mechanism in the original model with our SEC based Self-Attention and without introducing any other modules.
+As shown in Tab.
+2, we conduct experiments on image classification, object detection, insatance segmentation and semantic segmentation, the simple replacement of the attention mechanism yields significant advantages in both performance and efficiency.
+In addition to the supervised scenario, we also train the model with SimMIM [56] in the self-supervised scenario.
+As shown in Tab.
+3, SEC also performs exceptionally well in the self-supervised scenario.
+Image Classification.
+We compare our SECViT with numerous state-of-the-art models, the results are shown in Tab.4.
+We adopt the training strategy proposed in DeiT [46], with the only supervision is cross entropy loss.
+All of our models are trained from scratch for 300 epochs with the input resolution of $224\times 224$.
+SECViT consistently outperforms preceding models across all scales.
+Notably, SECViT-S attains a Top1-accuracy of 84.3% with a mere 27M parameters and 4.6G FLOPs.
+For larger models, SECViT-XL achieves a Top1-accuracy of 86.3% with 205M parameters and 36.4G FLOPs.
+The comparison of the models’ efficiency can be found in Appendix.
+Object Detection and Instance Segmentation.
+We utilize MMDetection [4] to implement Mask-RCNN [23], Cascade Mask R-CNN [2], and RetinaNet [36] to evaluate the performance of the SECViT.
 Tab.
@@
 The results show that SECViT performs better than its counterparts in all comparisons.
-For “3×+3$\times$+3 × +MS” schedule, SecViT outperforms the recent InternImage [50] for +2.2 box AP and +1.7 mask AP with the framework of Mask R-CNN.
-Regarding the Cascade Mask R-CNN, SecViT still has very good performance.
-As for “1×1$\times$1 ×” schedule, SecViT also demonstrate very impressive performance.
-Backbone Params (M) FLOPs (G) Mask R-CNN 1×1$\times$1 × Params (M) FLOPs (G) RetinaNet 1×1$\times$1 ×
-$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$ $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{b}_{S}$ $AP^{b}_{M}$ $AP^{b}_{L}$
-PVTv2-B1 [52] 33 243 41.8 54.3 45.9 38.8 61.2 41.6 23 225 41.2 61.9 43.9 25.4 44.5 54.3
-FAT-B2 [14] 33 215 45.2 67.9 49.0 41.3 64.6 44.0 23 196 44.0 65.2 47.2 27.5 47.7 58.8
-SecViT-T 34 221 47.7 69.5 52.6 42.9 66.4 46.1 24 202 45.6 66.8 48.9 28.8 49.8 60.7
-Swin-T [36] 48 267 43.7 66.6 47.7 39.8 63.3 42.7 38 248 41.7 63.1 44.3 27.0 45.3 54.7
-CMT-S [16] 45 249 44.6 66.8 48.9 40.7 63.9 43.4 44 231 44.3 65.5 47.5 27.1 48.3 59.1
-MPViT-S [28] 43 268 46.4 68.6 51.2 42.4 65.6 45.7 32 248 45.7 57.3 48.8 28.7 49.7 59.2
-STViT-S [24] 44 252 47.6 70.0 52.3 43.1 66.8 46.5
-FAT-B3 [14] 49 47.6 69.7 52.3 43.1 66.4 46.2 39 45.9 66.9 49.5 29.3 50.1 60.9
-BiFormer-S [62] 47.8 69.8 52.3 43.2 66.8 46.5 45.9 66.9 49.4 30.2 49.6 61.7
-SecViT-S 45 259 49.7 70.8 54.7 44.2 68.1 48.0 35 240 48.1 69.3 51.8 31.3 53.0 63.6
-Swin-S [36] 69 359 45.7 67.9 50.4 41.1 64.9 44.2 60 339 44.5 66.1 47.4 29.8 48.5 59.1
-ScalableViT-B [58] 95 349 46.8 68.7 51.5 42.5 65.8 45.9 85 330 45.8 67.3 49.2 29.9 49.5 61.0
-InternImage-S [50] 69 340 47.8 69.8 52.8 43.3 67.1 46.7
-CSWin-S [12] 54 342 47.9 70.1 52.6 43.2 67.1 46.2
-STViT-B [24] 70 359 49.7 71.7 54.7 44.8 68.9 48.7
-SecViT-B 73 368 51.2 72.7 56.4 45.4 69.7 49.0 63 349 48.9 70.2 52.7 32.0 53.6 64.6
-Swin-B [36] 107 496 46.9 69.2 51.6 42.3 66.0 45.5 98 477 45.0 66.4 48.3 28.4 49.1 60.6
-Focal-B [57] 110 533 47.8 70.2 52.5 43.2 67.3 46.5 101 514 46.3 68.0 49.8 31.7 50.4 60.8
-MPViT-B [28] 95 503 48.2 70.0 52.9 43.5 67.1 46.8 85 482 47.0 68.4 50.8 29.4 51.3 61.5
-CSwin-B [12] 97 526 48.7 70.4 53.9 43.9 67.8 47.3
-InternImage-B [50] 115 501 48.8 70.9 54.0 44.0 67.8 47.4
-SecViT-L 115 546 51.9 73.4 57.0 46.1 70.4 49.6 105 527 49.9 71.2 53.7 33.0 54.5 66.1
-Table 4: Comparison to other backbones using "1×1$\times$1 ד schedule on COCO.
-4.3 Semantic Segmentation
-Settings.
+Semantic Segmentation.
 We utilize Semantic FPN [29] and UperNet [55] to validate our SECViT’s performance, implementing these frameworks via MMSegmentation [7].
-We follow the PVT’s [51] training settings for Semantic FPN and train the model for 80k iterations.
-All models use an input resolution of $512\times 512$, and during testing, the image’s shorter side is resized to 512 pixels.
-Different from FPN, the UperNet is trained for 160K iterations and follows Swin’s [36] settings.
-We employ the AdamW optimizer with a weight decay of 0.01 and a 1500 iteration warm-up.
-Semantic FPN
-Backbone Params(M) FLOPs(G) mIoU(%)
-PVTv2-B1 [52] 18 34 42.5
-FAT-B2 [14] 17 32 45.4
-EdgeViT-S [39] 17 32 45.9
-SecViT-T 18 34 47.2
-DAT-T [54] 32 198 42.6
-CSWin-T [12] 26 202 48.2
-Shuted-S [44] 26 183 48.2
-FAT-B3 [14] 33 179 48.9
-SecViT-S 29 177 49.5
-DAT-S [54] 53 320 46.1
-RegionViT-B+ [5] 77 459 47.5
-UniFormer-B [29] 54 350 47.7
-CSWin-S [12] 39 271 49.2
-SecViT-B 57 288 50.7
-DAT-B [54] 92 481 47.0
-CrossFormer-L [53] 95 497 48.7
-CSWin-B [12] 81 464 49.9
-SecViT-L 99 471 52.0
-UperNet
-Backbone Params(M) FLOPs(G) mIoU(%)
-DAT-T [54] 60 957 45.5
-NAT-T [18] 58 934 47.1
-InternImage-T [50] 59 944 47.9
-MPViT-S [28] 52 943 48.3
-SMT-S [34] 50 935 49.2
-SecViT-S 56 933 50.6
-DAT-S [54] 81 1079 48.3
-SMT-B [34] 62 1004 49.6
-InterImage-S [50] 80 1017 50.2
-MPViT-B [28] 105 1186 50.3
-CSWin-S [12] 65 1027 50.4
-SecViT-B 84 1046 52.0
-Swin-B [36] 121 1188 48.1
-GC ViT-B [20] 125 1348 49.2
-DAT-B [54] 121 1212 49.4
-InternImage-B [50] 128 1185 50.8
-CSWin-B [12] 109 1222 51.1
-SecViT-L 130 1256 53.6
-Table 5: Comparison with the state-of-the-art on ADE20K.
-Results.
 The results of semantic segmentation can be found in the Tab.
-5.
-All the FLOPs are measured with the input resolution of $512\times 2048$, except the group of the SecViT-T, which are measured with the input resolution of $512\times 512$.
+7.
+All the FLOPs are measured with the input resolution of $512\times 2048$.
 SECViT achieves the best performance in all settings.
-Specifically, with the framework of Sementic FPN, our SecViT-S outperforms FAT-B3 for +0.6 mIoU.
-And SecViT-B surpasses CSWin-S by a large margin (+1.5 mIoU).
-For the framework of Upernet, SecViT-S surpasses the recent SMT-S by +1.4 mIoU.
-SecViT-B and SecViT-L also surpass all their counterparts.
+4.2 SEC for MLLM
+SEC can greatly facilitate the design of vision language connectors in MLLMs.
+First, we conduct a rigorous comparison between SEC and various baseline vision language connectors based on LLaVA-1.5.
+Then, we compare LLaVA-1.5+SEC with several popular contemporary MLLMs.
+Strict Comparison with Baselines.
+In Tab.
+8, we strictly compare various commonly used vision language connectors, including MLP, Resampler [1], Pooling, and EViT [35], which has achieved success in the design of ViT.
+Among these, MLP is the original design in LLaVA-1.5 [39], capable of achieving good results.
+However, it incurs significant computational cost due to the excessive vision tokens.
+To address this issue, some connectors attempt to use fewer vision tokens to accelerate LLaVA-1.5.
+Nonetheless, these adjustments inevitably lead to performance degradation.
+The results in Tab.
+8 show that using SEC can effectively accelerate the inference of LLaVA-1.5 without causing performance degradation, and can even improve the performance of LLaVA-1.5 to a certain extent.
+Comparison with Popular MLLMs.
+In Tab.
+9 and Tab.
+10, we compare LLaVA-1.5 equipped with SEC as a vision-language connector with other MLLMs.
+It is evident that SEC not only enhances the performance of MLLMs across various benchmarks but also significantly improves the efficiency of the models.
+This fully demonstrates the effectiveness of SEC in extracting visual information.
+4.3 Visualization of SEC
+To further understand the working mechanism of SEC, we visualize some clustering results for SECViT.
+As shown in Fig.
+5, the left side presents the clustering results of vision tokens at different stages of the model.
+From the clustering results, we analyze that in the shallow layers, the model distinguishes fine-grained features well, while in the deeper layers, it captures global semantic features effectively.
+The right side shows the Grad-CAM diagrams at different stages of the model, from which we can draw similar conclusions to the clustering results.
+More visualization results can be found in Appendix.
+Figure 5: Visualization for SEC.
 4.4 Ablation Study
-Comparison with Swin-Transformer.
-As mentioned earlier, based on the Swin-T architecture, we directly replaced WSA/SWSA with our SECA.
+In this section, we present some of the ablation study results for SEC, and more results can be found in the Appendix.
+Number of Vision Tokens in Each Clusters.
+The number of vision tokens has a significant impact on the performance and speed of the model.
+We thoroughly investigate the effect of the number of vision tokens on SECViT.
 As shown in Tab.
-6, this resulted in a significant performance improvement (classification accuracy increased by 2.3%).
-Comparison with CSWin-Transformer.
-As shown in Tab.
-6, based on the SecViT-T architecture, we replace SECA with CSWSA from CSwin-Transformer.
-This leads to a significant drop in classification performance (a decrease of 1.2%).
-This fully demonstrates the effectiveness of our proposed token clustering method.
-LCE.
-LCE is a simple depth-wise convolution module used to further enhance the model’s ability to model local information.
-We conduct ablation experiments on LCE and find that it contributes to a certain improvement in the model’s performance.
-The results are shown in Tab.
-6.
-Specifically, it boosts the model’s classification accuracy by +0.2.
-Model Params(M) FLOPs(G) Top1-acc(%)
-Swin-T 29 4.5 81.3
-SEC-Swin-T 29 4.9 83.6(+2.3)
-SecViT-T 15 2.5 82.6
-$\xrightarrow{}$ [12] 15 2.5 81.4(-1.2)
-w/o LCE 15 2.5 82.4
-w/o CPE 15 2.5 82.4
-w/o Conv Stem 14 2.1 82.3
-Table 6: Ablation Study.
-CPE.
-The CPE proposed in CPVT [7] is a plug-and-play, highly portable positional encoding method, widely adopted by many methods to provide positional information to the model.
-It consists of a residual block containing only a 3x3 depth-wise convolution.
-Ablation results on CPE in Tab.
-6 indicate that it offers some performance gains to the model (approximately +0.2 increase in classification accuracy).
-num FLOPs(G) Top1-acc(%)
-64, 16, 4, – 4.3 84.0
-64, 16, 2, – 4.3 84.1
-64, 8, 2, – 4.3 84.1
-32, 8, 2, – 4.4 84.2
-Table 7: Effect of number of clusters.
-Conv Stem.
-The Conv Stem is utilized in the early stages of the model to extract more refined local features.
-As indicated by the data in Tab.
-6, the Conv Stem enhances the performance of the model to some extent in classification task.
-Specifically, the Conv Stem improves the classification accuracy by +0.3.
-Effect of number of clusters.
-In Tab.
-7, we validate the impact of different cluster numbers chosen at each of the first three stages of the model on performance.
-For SecViT-S, we select 32, 8, and 2 clusters for the first three stages, respectively.
-This configuration achieves a classification accuracy of 84.2%.
-As we gradually increase the number of clusters at each stage, the model’s overall computation decreases, with a slight drop in accuracy.
-This demonstrates the robustness of Semantic Clustering with respect to the number of clusters.
-4.5 Visualization of Semantic Clustering
-Figure 4: Visualization results for SecViT-B.
-The first row shows the original images, the second row displays the clustering results of the first stage (32 clusters), the second row displays the clustering results of the second stage (8 clusters), and the third row shows the clustering results of the third stage (2 clusters).
-To further understand how our semantic equitable clustering works, we visualized the clustering results of some samples.
-As shown in Fig.
-4, we selected the first, second, and third stages of the pre-trained SecViT-B to observe the clustering results of different tokens.
-In the figure, patches with the same color belong to the same cluster.
-It is evident that in the first and second stages, our model tends to capture detailed information, while in the third stage, it tends to capture semantic information.
+11, the number of vision tokens in each cluster greatly influences the model’s performance.
+Specifically, in downstream dense prediction tasks, having too few tokens in each cluster leads to substantial performance degradation.
+When the number of tokens in each cluster is too large, the model’s performance does not see a significant improvement, but its speed decreases.
+V-T num Params (M) FLOPs (G) Throughput (imgs/s) Acc $AP^{b}$ $AP^{m}$ mIoU
+98 15 2.5 2004 82.7 47.8 43.0 47.2
+196 15 3.1 1722 83.0 (+0.3) 48.2 (+0.4) 43.4 (+0.4) 47.5 (+0.3)
+64 15 2.5 1946 82.7 (+0.0) 47.8 (+0.0) 42.8 (-0.2) 46.9 (-0.3)
+49 15 2.4 2102 82.6 (-0.1) 47.5 (-0.3) 42.7 (-0.3) 47.7 (-0.5)
+24 15 2.3 2186 82.0 (-0.7) 45.9 (-1.9) 40.6 (-2.4) 44.6 (-2.6)
+Table 11: Effect of the number of vision tokens in each cluster. “V-T num” means the number of vision tokens in each cluster. The experiments are conducted based on SECViT-T.
 5 Conclusion
-We propose a simple and straightforward clustering method for Vision Transformers—Semantic Equitable Clustering (SEC).
-This method assigns each token to a cluster by calculating the similarity between each token and a global token, then sorting them accordingly.
-Our clustering method takes into account the semantic information contained in the tokens, completes the clustering in a single step, and ensures an equal number of tokens in each cluster, facilitating efficient parallel processing on modern GPUs.
+We propose a simple and straightforward clustering method for vision tokens—Semantic Equitable Clustering (SEC).
+This method assigns each token to a cluster by calculating the similarity between each token and a global token, and completes the whole clustering process in only one step.
+Our clustering method takes into account the semantic information contained in the tokens, and ensures an equal number of tokens in each cluster, facilitating efficient parallel processing on modern GPUs.
 Based on Semantic Equitable Clustering, we designed SECViT, a versatile vision backbone that achieves impressive results across various vision tasks, including image classification, object detection, instance segmentation, and semantic segmentation.
-Supplementary Material
-Model Blocks Channels Heads Ratios Params(M) FLOPs(G)
-SecViT-T [2, 2, 9, 2] [64, 128, 256, 512] [2, 4, 8, 16] 3 15 2.5
-SecViT-S [3, 4, 18, 4] [64, 128, 256, 512] [2, 4, 8, 16] 3 27 4.4
-SecViT-B [4, 8, 26, 8] [80, 160, 320, 512] [5, 5, 10, 16] 3 55 9.7
-SecViT-L [4, 8, 26, 8] [112, 224, 448, 640] [7, 7, 14, 20] 3 96 18.1
-SEC-Swin-T [2, 2, 6, 2] [96, 192, 384, 768] [3, 6, 12, 24] 4 29 4.9
-SEC-Swin-S [2, 2, 18, 2] [96, 192, 384, 768] [3, 6, 12, 24] 4 50 9.3
-SEC-FasterViT-0 [2, 3, 6, 5] [64, 128, 256, 512] [–, –, 8, 16] 4 25 3.6
-SEC-FasterViT-1 [1, 3, 8, 5] [80, 160, 320, 640] [–, –, 8, 16] 4 41 5.6
-Table 8: Detailed Architectures of our models.
-Appendix A Architecture Details
+Besides, SEC can also be conveniently applied to multimodal large language models (MLLM) to serve as a vision language connector and benefits the model’s efficiency.
+Appendix A Experimental Details
+SECViT’s Architectures.
 SECViT’s architecture details are illustrated in Table 12.
@@
 $3\times 3$ DWConvs are adopted in CPE.
-While $5\times 5$ depth-wise convolutions are adopted for LCE.
-For SEC-Swin, we adhere to the design principles of the Swin-Transformer [36] without using additional structures such as CPE or Conv Stem.
-For SEC-FasterViT, we adhere to the design principles of FasterViT [19].
-We use convolutional for fast inference speed in the first two stages of the model and use our Semantic Equitable Clustering based Self-Attention (SECA) in the latter two stages.
 For all models, we set the number of clusters in the first three stages to 32, 8, and 2, respectively.
-Appendix B Experimental Settings
+Model Blocks Channels Heads Ratios Params(M) FLOPs(G)
+SECViT-T [2, 2, 9, 2] [64, 128, 256, 512] [2, 4, 8, 16] 3 15 2.5
+SECViT-S [4, 4, 18, 4] [64, 128, 256, 512] [2, 4, 8, 16] 3 27 4.6
+SECViT-B [4, 8, 26, 9] [80, 160, 320, 512] [2, 4, 8, 16] 3 57 9.8
+SECViT-L [4, 8, 26, 9] [112, 224, 448, 640] [4, 8, 14, 20] 3 101 18.2
+SECViT-XL [6, 12, 28, 12] [128, 256, 512, 1024] [4, 8, 16, 32] 3 205 36.4
+Table 12: Detailed Architectures of our models.
 ImageNet Image Classification.
@@
 The batch-size is set to 1024, respectively.
-We apply the same data augmentation and regularization used in DeiT [47], including RandAugment [9] (randm9-mstd0.5-inc1) , Mixup [60] (prob = 0.8), CutMix [59] (prob = 1.0), Random Erasing (prob = 0.25), and Exponential Moving Average (EMA) [41].
-The maximum rates of increasing stochastic depth [23] are set to 0.1/0.15/0.4/0.5 for SecViT-T/S/B/L.
+We apply the same data augmentation and regularization used in DeiT [46], including RandAugment [8] (randm9-mstd0.5-inc1) , Mixup [64] (prob = 0.8), CutMix [62] (prob = 1.0), Random Erasing (prob = 0.25).
+The maximum rates of increasing stochastic depth [24] are set to 0.1/0.15/0.4/0.5/0.65 for SECViT-T/S/B/L/XL.
 COCO Object Detection and Instance Segmentation.
 We apply RetinaNet [36], Mask-RCNN [23], and Cascaded Mask R-CNN [2] as the detection frameworks based on the MMDetection [4].
-All of our models are trained with “1 $\times$" (12 training epochs) and “3 $\times$ +MS" (36 training epochs with multi-scale training) settings.
-For the “1 $\times$" setting, images are resized to the shorter side of 800 pixels while the longer side is within 1333 pixels.
-For the “3 $\times$ +MS", multi-scale training strategy is used to randomly resize the shorter side of images between 480 to 800 pixels.
+All of our models are trained with “1 $\times$” (12 training epochs) and “3 $\times$ +MS” (36 training epochs with multi-scale training) settings.
+For the “1 $\times$” setting, images are resized to the shorter side of 800 pixels while the longer side is within 1333 pixels.
+For the “3 $\times$ +MS”, multi-scale training strategy is used to randomly resize the shorter side of images between 480 to 800 pixels.
 For both frameworks, we use the AdamW with the initial learning rate of 1e-4.
@@
 While for Mask-RCNN and Cascaded Mask R-CNN, we set it to 5e-2.
-ADE20K Semantic Segmentation.
+ADE20K Semantic Segmentaion.
 Based on MMSegmentation [7], we implement UperNet [55] and SemanticFPN [29] to evaluate our models’ performance on semantic segmentation.
@@
 For SemanticFPN, we also use the input resolution of $512\times 512$ but train the models for 80k iterations.
-Appendix C Limitations and Future Work
-While the proposed Semantic Equitable Clustering have demonstrated significant efficiency for clustering tokens and impressive performance on many vision tasks, there are still several limitations that need to be addressed in our future work.
-First, the computational constraints prevented us from experimenting with larger models and datasets such as ImageNet-21k.
-Exploring the potential of Semantic Equitable Clustering on large-scale datasets and larger model could provide further insights into its scalability and efficiency.
-In the future, we will strive to validate the performance of our SecViT on large datasets and with larger models.
-Appendix D Broader Impact Statement
-The development of the Semantic Equitable Clustering (SEC) has the potential to impact the field of computer vision by offering a efficient and semantic-related approach to group tokens in Vision Transformer.
-By clustering the tokens, SEC reduces computational load and improves the efficiency of vision models, which could lead to broader applications.
-The proposed SecViT is a general vision backbone that can be applied on different vision tasks, e.g., image classification, object detection, instance segmentation, and semantic segmentation.
-It has no direct negative social impact.
-Possible malicious uses of SecViT as a general-purpose backbone are beyond the scope of our study to discuss.
-Appendix E Code
-We provide the code of our Semantic Equitable Clustering base Self-Attention.
-1class SECA(nn.Module):
-2
-3 def __init__(self, embed_dim, num_heads, num_clusters):
-4 super().__init__()
-5 self.num_clusters = num_clusters
-6 self.embed_dim = embed_dim
-7 self.num_heads = num_heads
-8 self.head_dim = embed_dim // num_heads
-9 self.scaling = self.head_dim ** -0.5
-10 self.qkv = nn.Conv2d(embed_dim, embed_dim*3, 1, bias=True)
-11 self.out_proj = nn.Conv2d(embed_dim, embed_dim, 1, bias=True)
-12
-13 def sort_sim(self, q: torch.Tensor, k: torch.Tensor):
-14 ’’
-15 q: b n (h w) d
-16 k: b n (h w) d
-17 ’’
-18 cls_token = q.mean(2, keepdim=True) # (b n 1 d)
-19 cls_token = F.normalize(cls_token, dim=-1)
-20 k = F.normalize(k, dim=-1)
-21 cos_sim = cls_token @ k.transpose(-1, -2) # (b n 1 (h w))
-22 sort_indices = torch.argsort(cos_sim, dim=-1)
-23 return sort_indices.squeeze(2) # (b n (h w))
-24
-25 def sort_tensor(self, indices: torch.Tensor, target: torch.Tensor):
-26 ’’
-27 indices: (b n l)
-28 target: (b n l d)
-29 ’’
-30 sorted_tensor = torch.gather(target, dim=-2, index=indices.unsqueeze(-1).expand_as(target))
-31 return sorted_tensor
-32
-33 def restore_tensor(self, indices: torch.Tensor, target: torch.Tensor):
-34 ’’
-35 indices: (b n l)
-36 target: (b n l d)
-37 ’’
-38 restored_tensor = torch.gather(target, dim=-2, index=indices.sort(dim=-1).indices.unsqueeze(-1).expand_as(target))
-39 return restored_tensor
-40
-41 def forward(self, x: torch.Tensor):
-42 ’’
-43 x: (b c h w)
-44 ’’
-45 bsz, _, h, w = x.size()
-46 l = h * w
-47 assert l % self.num_clusters == 0
-48 qkv = self.qkv(x) # (b 3*c h w)
-49
-50 q, k, v = rearrange(qkv, ’b (m n d) h w -> m b n (h w) d’, m=3, n=self.num_heads)
-51 # q, k, v -> b n (h w) d
-52 k = k * self.scaling
-53
-54 if self.num_clusters > 1:
-55 indices = self.sort_sim(k, k) # (b n (h w))
-56 q = self.sort_tensor(indices, q) # (b n (h w) d)
-57 k = self.sort_tensor(indices, k) # (b n (h w) d)
-58 v = self.sort_tensor(indices, v) # (b n (h w) d)
-59
-60 q_l = rearrange(q, ’b n (c l) d -> b n c l d’, c=self.num_clusters)
-61 k_l = rearrange(k, ’b n (c l) d -> b n c l d’, c=self.num_clusters)
-62 v = rearrange(v, ’b n (c l) d -> b n c l d’, c=self.num_clusters)
-63 attn = torch.softmax(q_l @ k_l.transpose(-1, -2), dim=-1) # (b n c l l)
-64 v = attn @ v # (b n c l d)
-65
-66 v = rearrange(v, ’b n c l d -> b n (c l) d’)
-67 v = self.restore_tensor(indices, v) # (b n (h w) d)
-68 v = rearrange(v, ’b n (h w) d -> b (n d) h w’, h=h, w=w)
-69 else:
-70 attn = torch.softmax(q @ k.transpose(-1, -2), dim=-1) # (b n (h w) (h w))
-71 v = attn @ v # (b n (h w) d)
-72 v = rearrange(v, ’b n (h w) d -> b (n d) h w’, h=h, w=w)
-73 return self.out_proj(res)
+Appendix B More experimental Results
+Efficiency Comparison.
+In Tab. 13, we compare the inference efficiency of various models in detail. From this, we can see that the ViT based on SEC demonstrates the best performance-speed tradeoff.
+Model Params(M) FLOPs(G) Throughput(imgs/s) Top1-Acc(%)
+DeiT-S [46] 22 4.6 3204 79.8
+EViT-DeiT-S (keeprate=0.9) [35] 22 4.0 3428 79.8
+SEC-DeiT-S (num_cluster=4) 22 4.1 3412 80.5
+DeiT-B [46] 86 17.6 1502 81.8
+SEC-DeiT-B 86 14.8 1682 82.4
+PVTv2-b1 [49] 13 2.1 2204 78.7
+TCFormer-light [63] 14 3.8 417 79.4
+MPViT-XS [31] 11 2.9 1496 80.9
+BiFormer-T [66] 13 2.2 1634 81.4
+CMT-XS [18] 15 1.5 1476 81.8
+GC-ViT-XT [22] 20 2.6 1308 82.0
+SMT-T [37] 12 2.4 638 82.2
+RMT-T [15] 14 2.5 1438 82.4
+SECViT-T 15 2.5 2004 82.7
+Swin-T [41] 29 4.5 1723 81.3
+PS-ViT-B14 [61] 21 5.4 1986 81.7
+DVT-T2T-ViT-19 [52] 39 6.2 1268 81.9
+SGFormer-S [17] 23 4.8 952 83.2
+CMT-S [18] 25 4.0 846 83.5
+CSwin-S [12] 35 6.9 972 83.6
+SMT-S [37] 20 4.8 356 83.7
+BiFormer-S [66] 26 4.5 766 83.8
+SEC-Swin-T 29 4.8 1482 83.8
+SECViT-S 27 4.6 998 84.3
+Swin-S [41] 50 8.8 1006 83.0
+SGFormer-M [17] 39 7.5 598 84.1
+SMT-B [37] 32 7.7 237 84.3
+BiFormer-B [66] 57 9.8 498 84.3
+MaxViT-S [47] 69 11.7 546 84.5
+CMT-B [18] 46 9.3 447 84.5
+iFormer-B [44] 48 9.4 688 84.6
+RMT-B [15] 54 9.7 430 85.0
+SEC-Swin-S 50 9.2 804 85.0
+SECViT-B 57 9.8 504 85.2
+Swin-B [41] 88 15.5 768 83.5
+CSWin-B [12] 78 15.0 660 84.2
+SMT-L [37] 80 17.7 158 84.6
+SGFormer-B [17] 78 15.6 388 84.7
+iFormer-L [44] 87 14.0 410 84.8
+MaxViT-B [47] 120 23.4 306 84.9
+SEC-Swin-B 88 16.2 696 85.3
+SECViT-L 101 18.2 398 85.7
+Table 13: Comparison of models’ efficiency. Throughputs are measured on a single A100 with the batch size of 64.
+Number of Vision Tokens Outputs by SEC.
+MLLM is quite sensitive to the number of vision tokens.
+We conduct a detailed exploration based on LLaVA-1.5 regarding the number of vision tokens output by SEC, as shown in Tab.
+14.
+The first row represents the speed and performance of the original LLaVA-1.5 without using SEC.
+Compared to LLaVA-1.5, employing SEC effectively reduces the number of vision tokens and improves inference efficiency.
+As the number of vision tokens decreases, the model’s performance shows a slight decline, but its efficiency is further enhanced.
+V-T num Time Speed TextVQA GQA VQAv2 POPE MM-Vet
+576+1 194s 1.0×1.0$\times$1.0 × 58.2 62.0 78.5 86.1 31.1
+288+1 126s 1.5×1.5$\times$1.5 × 60.1(+1.9) 63.5(+1.5) 78.9(+0.4) 87.7(+1.6) 33.2(+2.1)
+256+1 116s 1.7×1.7$\times$1.7 × 59.6(+1.4) 63.2(+0.3) 78.6(+0.1) 87.1(+1.0) 32.7(+1.6)
+192+1 102s 1.9×1.9$\times$1.9 × 57.7(-0.5) 62.7(+0.7) 78.4(-0.1) 86.7(+0.6) 32.1(+1.0)
+144+1 94s 2.1×2.1$\times$2.1 × 56.8(-1.4) 62.0(+0.0) 78.0(-0.5) 86.1(+0.0) 31.7(+0.6)
+Table 14: Effect of the number of vision tokens outputs by SEC. “V-T num” means the number of vision tokens output by SEC. The experiments are conducted based on LLaVA-1.5 [39].
+Different Methods for Merging Vision Tokens.
+For MLLM, SEC uses an interleaved merge token approach to reduce the number of vision tokens.
+Conversely, we also explore a sequential merge token method to achieve a similar reduction.
+The comparison of these two methods is shown in Tab.
+15.
+The direct sequential merge token approach may result in the loss of critical visual information, significantly degrading the model’s performance.
+Method V-T num Time Speed TextVQA GQA VQAv2 POPE MM-Vet
+Interleaved 288+1 14h 1.5×1.5$\times$1.5 × 60.1 63.5 78.9 87.7 33.2
+Sequential 288+1 14h 1.5×1.5$\times$1.5 × 52.8(-7.3) 57.1(-6.2) 75.7(-3.2) 81.7(-6.0) 27.6(-5.6)
+Interleaved 144+1 10h 2.1×2.1$\times$2.1 × 56.8 62.0 78.0 86.1 31.7
+Sequential 144+1 10h 2.1×2.1$\times$2.1 × 47.2(-9.6) 53.6(-8.4) 71.7(-6.3) 80.0(-6.1) 22.3(-9.6)
+Table 15: Different methods for merging vision tokens.
+Appendix C More Clustering Results for Complex Scenes
+To further illustrate the mechanism of SEC, we visualize more images in complex scenes and their clustering results, as shown in Fig.
+6.
+Specifically, we visualize the clustering results of the first three stages of SECViT.
+The results further demonstrate that SEC can better learn fine-grained representations in the shallow layers of the model and semantic representations in the deeper layers.
+Figure 6: Visualization results for complex scenes.

### v3

@@
 Attention is calculated only within the same cluster.
-Models Params (M) FLOPs (G) Throughput (imgs/s) Acc $AP^{b}$ $AP^{m}$ mIoU
-DeiT-S 22 4.6 3204 79.8 44.5 40.1 43.0
-EViT-DeiT-S (keeprate=0.9) 22 4.0 3428 79.8 not suit not suit not suit
-SEC-DeiT-S (num_cluster=2) 22 4.3 3226 80.6 (+0.8) 47.9 (+3.4) 43.1 (+3.0) 48.0 (+5.0)
-SEC-DeiT-S (num_cluster=4) 22 4.1 3412 80.5 (+0.7) 47.7 (+3.2) 42.7 (+2.6) 47.5 (+4.5)
-SEC-DeiT-S (num_cluster=8) 22 3.9 3528 80.1 (+0.3) 46.7 (+2.2) 42.0 (+1.9) 46.4 (+3.4)
-Table 1: Comparison of EViT [35] and SEC. Inference speed are measured on the A100 GPU.
-Different adaptability to vision models/MLLMs.
-In pruning methods like EViT, during the model’s forward pass, the number of tokens gradually decreases.
-Although reducing tokens does not impact classification tasks, it prevents the feature map from being restored to its original shape.
-This makes it difficult for EViT to be directly used with classic frameworks like SemanticFPN for downstream dense prediction tasks .
-In SEC, we simply group the tokens without changing their quantity, thereby preserving the integrity of the feature map.
-This ensures that SEC can be easily applied to downstream tasks.
-In Tab.
-1, we use DeiT [46] as the baseline.
-Without introducing any other tricks, just by applying SEC to DeiT, we form SEC-DeiT and compare it with EViT-DeiT.
-We conducted image classification, object detection, instance segmentation and semantic segmentation based on SEC-DeiT-S.
-SEC not only accelerates the model but also enhances its performance.
-Beyond the vision tasks, we also compare the EViT with our SEC on MLLM, details can be found in the Tab.
-8.
 4 Experiments
@@
 SECViT-L 101 18.2 85.7
-XL model $\sim 35.0$G ConvNeXt-L [42] 198 34.4 84.3
-CoAtNet-3 [10] 168 34.7 84.5
-MaxViT-L [47] 212 43.9 85.1
-GC ViT-L [22] 201 32.6 85.7
-SECViT-XL 205 36.4 86.3
 Table 3: Comparison with the state-of-the-art on ImageNet-1K classification.
@@
 Cascade Mask R-CNN 3×3$\times$3 ×+MS
-NAT-T [21] 85 737 51.4 70.0 55.9 44.5 67.6 47.9
 GC-ViT-T [20] 85 770 51.6 70.4 56.1 44.6 67.8 48.3
@@
 MogaNet-L [31] 113 1176 50.9
-CSWin-B [12] 81 464 49.9 109 1222 51.1
 MLLA-B [18] 128 1183 51.9
@@
 As shown in Tab.
-2, we conduct experiments on image classification, object detection, insatance segmentation and semantic segmentation, the simple replacement of the attention mechanism yields significant advantages in both performance and efficiency.
+1, we conduct experiments on image classification, object detection, insatance segmentation and semantic segmentation, the simple replacement of the attention mechanism yields significant advantages in both performance and efficiency.
 In addition to the supervised scenario, we also train the model with SimMIM [53] in the self-supervised scenario.
 As shown in Tab.
-3, SEC also performs exceptionally well in the self-supervised scenario.
+2, SEC also performs exceptionally well in the self-supervised scenario.
 Image Classification.
@@
 Notably, SECViT-S attains a Top1-accuracy of 84.3% with a mere 27M parameters and 4.6G FLOPs.
-For larger models, SECViT-XL achieves a Top1-accuracy of 86.3% with 205M parameters and 36.4G FLOPs.
 The comparison of the models’ efficiency can be found in Appendix.
@@
 The results of semantic segmentation can be found in the Tab.
-7.
+6.
 All the FLOPs are measured with the input resolution of $512\times 2048$.
@@
 In Tab.
-8, we strictly compare various commonly used vision language connectors, including MLP, Resampler [1], Pooling, and EViT [35], which has achieved success in the design of ViT.
+7, we strictly compare various commonly used vision language connectors, including MLP, Resampler [1], Pooling, and EViT [32], which has achieved success in the design of ViT.
 Among these, MLP is the original design in LLaVA-1.5 [36], capable of achieving good results.
@@
 8 and Tab.
-10, we compare LLaVA-1.5 equipped with SEC as a vision-language connector with other MLLMs.
+9, we compare LLaVA-1.5 equipped with SEC as a vision-language connector with other MLLMs.
 It is evident that SEC not only enhances the performance of MLLMs across various benchmarks but also significantly improves the efficiency of the models.
 This fully demonstrates the effectiveness of SEC in extracting visual information.
-4.3 Visualization of SEC
-To further understand the working mechanism of SEC, we visualize some clustering results for SECViT.
-As shown in Fig.
-5, the left side presents the clustering results of vision tokens at different stages of the model.
-From the clustering results, we analyze that in the shallow layers, the model distinguishes fine-grained features well, while in the deeper layers, it captures global semantic features effectively.
-The right side shows the Grad-CAM diagrams at different stages of the model, from which we can draw similar conclusions to the clustering results.
-More visualization results can be found in Appendix.
-Figure 5: Visualization for SEC.
 4.3 Ablation Study
@@
 As shown in Tab.
-11, the number of vision tokens in each cluster greatly influences the model’s performance.
+10, the number of vision tokens in each cluster greatly influences the model’s performance.
 Specifically, in downstream dense prediction tasks, having too few tokens in each cluster leads to substantial performance degradation.
@@
 Table 10: Effect of the number of vision tokens in each cluster. “V-T num” means the number of vision tokens in each cluster. The experiments are conducted based on SECViT-T.
+4.4 Visualization of SEC
+To further understand the working mechanism of SEC, we visualize some clustering results for SECViT.
+As shown in Fig.
+5, the left side presents the clustering results of vision tokens at different stages of the model.
+From the clustering results, we analyze that in the shallow layers, the model distinguishes fine-grained features well, while in the deeper layers, it captures global semantic features effectively.
+The right side shows the Grad-CAM diagrams at different stages of the model, from which we can draw similar conclusions to the clustering results.
+More visualization results can be found in Appendix.
+Figure 5: Visualization for SEC.
+Number of Vision Tokens Outputs by SEC.
+MLLM is quite sensitive to the number of vision tokens.
+We conduct a detailed exploration based on LLaVA-1.5 regarding the number of vision tokens output by SEC, as shown in Tab.
+11.
+The first row represents the speed and performance of the original LLaVA-1.5 without using SEC.
+Compared to LLaVA-1.5, employing SEC effectively reduces the number of vision tokens and improves training efficiency.
+As the number of vision tokens decreases, the model’s performance shows a slight decline, but its efficiency is further enhanced.
+V-T num Time Speed TextVQA GQA VQAv2 POPE MM-Vet
+576+1 21h 1.0×1.0$\times$1.0 × 58.2 62.0 78.5 86.1 31.1
+288+1 14h 1.5×1.5$\times$1.5 × 60.1(+1.9) 63.5(+1.5) 78.9(+0.4) 87.7(+1.6) 33.2(+2.1)
+256+1 13h 1.6×1.6$\times$1.6 × 59.6(+1.4) 63.2(+0.3) 78.6(+0.1) 87.1(+1.0) 32.7(+1.6)
+192+1 11h 1.9×1.9$\times$1.9 × 57.7(-0.5) 62.7(+0.7) 78.4(-0.1) 86.7(+0.6) 32.1(+1.0)
+144+1 10h 2.1×2.1$\times$2.1 × 56.8(-1.4) 62.0(+0.0) 78.0(-0.5) 86.1(+0.0) 31.7(+0.6)
+Table 11: Effect of the number of vision tokens outputs by SEC. “V-T num” means the number of vision tokens output by SEC. The experiments are conducted based on LLaVA-1.5 [36].
 5 Conclusion
@@
 Besides, SEC can also be conveniently applied to multimodal large language models (MLLM) to serve as a vision language connector and benefits the model’s efficiency.
+6 Acknowledgements
+This work is partially funded by Beijing Natural Science Foundation (4252054), Youth Innovation Promotion Association CAS(Grant No.2022132), Beijing Nova Program(20230484276), and CCF-Kuaishou Large Model Explorer Fund (NO.
+CCF-KuaiShou 2024005).
+[…]
+Supplementary Material
 Appendix A Experimental Details
@@
 Table 12: Detailed Architectures of our models.
-ImageNet Image Classification.
-We adopt the training strategy proposed in DeiT [46], with the only supervision is cross entropy loss.
-All of our models are trained from scratch for 300 epochs with the input resolution of $224\times 224$.
-The AdamW is used with a cosine decay learning rate scheduler and 5 epochs of linear warm-up.
-The batch-size is set to 1024, respectively.
-We apply the same data augmentation and regularization used in DeiT [46], including RandAugment [8] (randm9-mstd0.5-inc1) , Mixup [64] (prob = 0.8), CutMix [62] (prob = 1.0), Random Erasing (prob = 0.25).
-The maximum rates of increasing stochastic depth [24] are set to 0.1/0.15/0.4/0.5/0.65 for SECViT-T/S/B/L/XL.
-COCO Object Detection and Instance Segmentation.
-We apply RetinaNet [36], Mask-RCNN [23], and Cascaded Mask R-CNN [2] as the detection frameworks based on the MMDetection [4].
-All of our models are trained with “1 $\times$” (12 training epochs) and “3 $\times$ +MS” (36 training epochs with multi-scale training) settings.
-For the “1 $\times$” setting, images are resized to the shorter side of 800 pixels while the longer side is within 1333 pixels.
-For the “3 $\times$ +MS”, multi-scale training strategy is used to randomly resize the shorter side of images between 480 to 800 pixels.
-For both frameworks, we use the AdamW with the initial learning rate of 1e-4.
-For RetinaNet, we set the weight decay to 1e-4.
-While for Mask-RCNN and Cascaded Mask R-CNN, we set it to 5e-2.
-ADE20K Semantic Segmentaion.
-Based on MMSegmentation [7], we implement UperNet [55] and SemanticFPN [29] to evaluate our models’ performance on semantic segmentation.
-For UperNet, we follow the previous setting of Swin-Transformer [41] and train the model for 160k iterations with the input size of $512\times 512$.
-For SemanticFPN, we also use the input resolution of $512\times 512$ but train the models for 80k iterations.
 Appendix B More experimental Results
@@
 Table 13: Comparison of models’ efficiency. Throughputs are measured on a single A100 with the batch size of 64.
-Number of Vision Tokens Outputs by SEC.
-MLLM is quite sensitive to the number of vision tokens.
-We conduct a detailed exploration based on LLaVA-1.5 regarding the number of vision tokens output by SEC, as shown in Tab.
-14.
-The first row represents the speed and performance of the original LLaVA-1.5 without using SEC.
-Compared to LLaVA-1.5, employing SEC effectively reduces the number of vision tokens and improves inference efficiency.
-As the number of vision tokens decreases, the model’s performance shows a slight decline, but its efficiency is further enhanced.
-V-T num Time Speed TextVQA GQA VQAv2 POPE MM-Vet
-576+1 194s 1.0×1.0$\times$1.0 × 58.2 62.0 78.5 86.1 31.1
-288+1 126s 1.5×1.5$\times$1.5 × 60.1(+1.9) 63.5(+1.5) 78.9(+0.4) 87.7(+1.6) 33.2(+2.1)
-256+1 116s 1.7×1.7$\times$1.7 × 59.6(+1.4) 63.2(+0.3) 78.6(+0.1) 87.1(+1.0) 32.7(+1.6)
-192+1 102s 1.9×1.9$\times$1.9 × 57.7(-0.5) 62.7(+0.7) 78.4(-0.1) 86.7(+0.6) 32.1(+1.0)
-144+1 94s 2.1×2.1$\times$2.1 × 56.8(-1.4) 62.0(+0.0) 78.0(-0.5) 86.1(+0.0) 31.7(+0.6)
-Table 14: Effect of the number of vision tokens outputs by SEC. “V-T num” means the number of vision tokens output by SEC. The experiments are conducted based on LLaVA-1.5 [39].
 Different Methods for Merging Vision Tokens.
@@
 The comparison of these two methods is shown in Tab.
-15.
+14.
 The direct sequential merge token approach may result in the loss of critical visual information, significantly degrading the model’s performance.